### v1

# SemEval-2025 Task 11: Bridging the Gap in Text-Based Emotion Detection

Abstract
We present our shared task on text-based emotion detection, covering more than 30 languages from seven distinct language families. These languages are predominantly low-resource and spoken across various continents. The data instances are multi-labeled into six emotional classes, with additional datasets in 11 languages annotated for emotion intensity. Participants were asked to predict labels in three tracks: (a) emotion labels in monolingual settings, (b) emotion intensity scores, and (c) emotion labels in cross-lingual settings. The task attracted over 700 participants. We received final submissions from more than 200 teams and 93 system description papers. We report baseline results, as well as findings on the best-performing systems, the most common approaches, and the most effective methods across various tracks and languages. The datasets for this task are publicly available.  [^1] The dataset is available at SemEval2024-task 11.
[…]
SemEval-2025 Task 11: Bridging the Gap in Text-Based Emotion Detection
Shamsuddeen Hassan Muhammad1,2thanks: Equal contribution, Nedjma Ousidhoum3∗, Idris Abdulmumin4, Seid Muhie Yimam5, Jan Philip Wahle6, Terry Ruas6, Meriem Beloucif7, Christine De Kock8, Tadesse Destaw Belay9,10, Ibrahim Said Ahmad11, Nirmal Surange12, Daniela Teodorescu13, David Ifeoluwa Adelani14,15,16, Alham Fikri Aji17, Felermino Ali18, Vladimir Araujo19, Abinew Ali Ayele5,20, Oana Ignat21, Alexander Panchenko22,23, Yi Zhou3, Saif M. Mohammad24 1Imperial College London, 2Bayero University Kano, 3Cardiff University, 4DSFSI, University of Pretoria, 5University of Hamburg, 6University of Göttingen, 7Uppsala University, 8University of Melbourne, 9Instituto Politécnico Nacional, 10Wollo University, 11Northeastern University,12IIIT Hyderabad, 13University of Alberta, 14MILA, 15McGill University, 16Canada CIFAR AI Chair,17MBZUAI,18LIACC, FEUP, University of Porto, 19Sailplane AI, 20Bahir Dar University, 21Santa Clara University, 22Skoltech, 23AIRI, 24National Research Council Canada Contact: s.muhammad@imperial.ac.uk, OusidhoumN@cardiff.ac.uk
1 Introduction
People use language in diverse and sophisticated ways to express emotions across languages and cultures Wiebe et al. (2005); Mohammad and Kiritchenko (2018); Mohammad et al. (2018a). Emotions are also perceived subjectively, even within the same culture or social group. Recognising these emotions is central to language technologies and NLP applications in healthcare, digital humanities, dialogue systems, and beyond Mohammad et al. (2018b); Saffar et al. (2023). In this work, we use emotion recognition to refer to perceived emotions—i.e., the emotion most people believe the speaker might have felt based on a sentence or short text snippet.
Figure 1: Languages in the three tracks (A, B, and C) of SemEval 2025 Task 11.
Despite the linguistic diversity of regions such as Africa and Asia, which together account for more than 4,000 languages [^2] https://www.ethnologue.com/insights/how-many-languages, few emotion recognition resources exist for these languages. Prior SemEval shared tasks on emotion recognition have primarily focused on high-resource languages such as English, Spanish, and Arabic Strapparava and Mihalcea (2007); Mohammad et al. (2018a); Chatterjee et al. (2019). In this task, we provide participants with new datasets covering more than 30 languages from seven distinct language families, spoken across Africa, Asia, Latin America, North America, and Europe Muhammad et al. (2025). Our manually annotated emotion recognition datasets, curated in collaboration with local communities, consist of over 100,000 multi-labeled instances drawn from diverse sources, including speeches, social media, news, literature, and reviews. Each instance is labeled by fluent speakers and annotated with six emotion classes: joy, sadness, anger, fear, surprise, disgust, and neutral. Additionally, eleven datasets include four emotional intensity levels ranging from 0 to 3 (i.e., absence of emotion to high intensity).
The task consists of three tracks: (a) multi-label emotion detection, (b) emotion intensity detection, and (c) cross-lingual emotion detection. The languages for each track are listed in Figure 1. Each team could submit results for one, two, or all three tracks in one or more languages. Our official evaluation metrics were the average F-score for Tracks A and C and the Pearson correlation coefficient for Track B, which measures how well system-predicted intensity scores align with human judgments.
Our task attracted over 700 participants, with 220 final submissions and 93 teams submitting system description papers. Track A (multi-label emotion detection) received the most submissions (114), followed by Track C (cross-lingual emotion detection) with 51, and Track B (emotion intensity detection) with 32. Most teams participated in multiple languages, averaging 11 languages per team. Our task was the most popular competition on CodaBench in 2024.  [^3] Codabench Newsletter 2024 All task details and resources are available on the task’s GitHub page. [^4] https://github.com/emotion-analysis-project/SemEval2025-Task11
2 Related Work
NLP work on emotion detection is predominantly Western-centric, with a few exceptions for languages other than English (e.g., Italian (Bianchi et al., 2021), Romanian (Ciobotaru et al., 2022), Indonesian (Saputri et al., 2018), and Bengali (Iqbal et al., 2022)). While multilingual datasets (e.g., (Öhman et al., 2020) and XLM-EMO (Bianchi et al., 2022)) exist, they do not fully capture cultural nuances in emotional expressions due to their reliance on translated data (e.g., XLM-EMO), as emotions are highly contextualized and culture-specific (Havaldar et al., 2023; Mohamed et al., 2024; Hershcovich et al., 2022). Furthermore, most datasets are single-labeled, and to the best of our knowledge, there are no multilingual resources that capture overlapping emotions and their intensity across various languages.
Additionally, most prior emotion recognition shared tasks have focused on high-resource languages such as English, Spanish, German, and Arabic Strapparava and Mihalcea (2007); Mohammad et al. (2018b); Chatterjee et al. (2019). In contrast, this shared task covers more than 30 languages, including many low-resource languages.
3 Data
Language Train Dev Test Total
Afrikaans (afr) 2,107 98 1,065 3,270
Amharic (amh) 3,549 592 1,774 5,915
Algerian Arabic (arq) 901 100 902 1,903
Moroccan Arabic (ary) 1,608 267 812 2,687
Chinese (chn) 2,642 200 2,642 5,484
German (deu) 2,603 200 2,604 5,407
English (eng) 2,768 116 2,767 5,651
Latin American Spanish (esp) 1,996 184 1,695 3,875
Hausa (hau) 2,145 356 1,080 3,581
Hindi (hin) 2,556 100 1,010 3,666
Igbo (ibo) 2,880 479 1,444 4,803
Indonesian (ind) 156 851 1,007
Javanese (jav) 151 837 988
Kinyarwanda (kin) 2,451 407 1,231 4,089
Marathi (mar) 2,415 100 1,000 3,515
Nigerian-Pidgin (pcm) 3,728 620 1,870 6,218
Oromo (orm) 3,442 575 1,721 5,738
Portuguese (Brazilian; ptbr) 2,226 200 2,226 4,652
Portuguese (Mozambican; ptmz) 1,546 257 776 2,579
Romanian (ron) 1,241 123 1,119 2,483
Russian (rus) 2,679 199 1,000 3,878
Somali (som) 3,392 566 1,696 5,654
Sundanese (sun) 924 199 926 2,049
Swahili (swa) 3,307 551 1,656 5,514
Swedish (swe) 1,187 200 1,188 2,575
Tatar (tat) 1,000 200 1,000 2,200
Tigrinya (tir) 3,681 614 1,840 6,135
Ukrainian (ukr) 2,466 249 2,234 4,949
Emakhuwa (vmw) 1,551 258 777 2,586
isiXhosa (xho) 682 1,594 2,276
Yoruba (yor) 2,992 497 1,500 4,989
isiZulu (zul) 875 2,047 2,922
Table 1: Languages and data split sizes. Datasets with no training splits (-) were only used in Track C (crosslingual) only.
3.1 Data Collection
As our task includes more than 30 different datasets, curated and annotated by fluent speakers, we selected data sources based on: 1) the availability of textual data potentially rich in emotions, and 2) access to annotators. Since finding suitable data is challenging when resources are limited, we typically combine sources. The main textual sources used to build our dataset collection are:
  • Social media posts: Data collected from various platforms, including Reddit (e.g., eng, deu), YouTube (e.g., esp, ind, jav, sun, tir), Twitter (e.g., amh, hau), and Weibo (e.g., chn).
  • Personal narratives, talks, speeches: Anonymised sentences from personal diary posts. We use these in eng, deu, and ptbr, mainly from subreddits such as IAmI. Similarly, the afr dataset includes sentences from speeches and talks.
  • Literary texts: We translated the novel *La Grande Maison* (The Big House) by Mohammed Dib  [^5] https://en.wikipedia.org/wiki/La_Grande_Maison from French to Algerian Arabic (arq) and manually post-processed the translation to generate sentences to be annotated by native speakers.
  • News data: Although we prefer emotionally rich social media data from different platforms, when such data is scarce, we annotate news data and headlines in some African languages (e.g., yor, hau, and vmw).
  • Human-written and machine-generated data: We created a dataset from scratch for Hindi (hin) and Marathi (mar). We ask annotators to generate emotive sentences on a given topic (e.g., family), automatically translate a small section of the Hindi dataset to Marathi, and have native speakers manually fix translation errors. Finally, we augment both datasets with a few hundred quality-approved instances generated by ChatGPT.
3.2 Data Annotation
We ask the annotators to select all the emotions that apply to a given text. The set of perceived emotion labels includes: anger, sadness, fear, disgust, joy, surprise, and neutral (if no emotion is present). The annotators further rate the selected emotion(s) on a four-point intensity scale: 0 (no emotion), 1 (low intensity), 2 (moderate intensity), and 3 (high intensity). We provide the definitions of the categories, annotation guidelines, and more details in Muhammad et al. (2025). We expected some level of disagreement, as emotions are complex, subtle, and perceived differently, even by people within the same culture, especially in the absence of full context. Hence, the final emotion labels were determined based on the emotions and associated intensity values selected by the annotators. Specifically, the given emotion is considered present if:
  1. 1.
    At least two annotators select a label with an intensity value of 1, 2, or 3 (low, medium, or high, respectively).
  2. 2.
    The average score exceeds a predefined threshold $T$. We set $T$ to $0.5$.
Once the labels for perceived emotions are assigned, the final intensity scores for Track B were determined by averaging the selected intensity values and rounding up to the nearest whole number. We only assign intensity scores for datasets where most instances were annotated by at least five annotators to ensure robustness. Table 1 shows the total number of instances in each dataset and the number of instances in the training, development, and test splits for all the languages.
3.3 Annotators’ Reliability
We report the reliability of the annotation using the Split-Half Class Match Percentage (SHCMP; Mohammad, 2024) as described in Muhammad et al. (2025). SHCMP extends the concept of Split-Half Reliability (SHR), traditionally used for continuous scores (Kiritchenko and Mohammad, 2016), to discrete categories like ours (i.e., intensity scores per emotion). Overall, the scores vary from 60% to more than 90%, indicating that our datasets are of high quality.
4 Task Description
Participants were given text snippets and asked to determine the emotions that people may attribute to the speaker based on a sentence or short text snippet uttered by the speaker. The task consists of three tracks, and participants could participate in one or more of these tracks.
4.1 Tracks
Track A: Multi-label Emotion Detection
Participants were asked to predict the perceived emotion(s) of the speaker and label each text snippet based on the presence (1) or absence (0) of the following emotions: joy, sadness, fear, anger, surprise, and disgust.
Track B: Emotion Intensity Detection
Given a text and the six emotion classes (i.e., joy, sadness, fear, anger, surprise, and disgust), participants had to predict whether the intensity of each possible emotion was equal to 0 (no emotion), 1 (low), 2 (medium), or 3 (high).
Track C: Cross-lingual Emotion Detection
Similar to Track A, participants had to predict the presence or absence of each perceived emotion, but without using any training data in the target language. Instead, participants were allowed to use labeled dataset(s) from at least one other language. For instance, one could use German data for training when testing on English.
4.2 Task Organisation
Figure 2: The official affiliations of some of our participants. The list includes 33 countries: Argentina, Australia, Bangladesh, Brazil, Canada, Colombia, Egypt, Ethiopia, Finland, Germany, Greece, India, Indonesia, Iran, Japan, Jordan, Mexico, Morocco, the Netherlands, Nigeria, Pakistan, Poland, Romania, Russia, South Africa, Spain, Sweden, Taiwan, the UAE, the UK, the USA, and Vietnam.
We used Codabench as the competition platform and released pilot datasets before the start of the shared task to help participants better understand the task (i.e., the datasets, the languages involved, and the labels). We provided participants with a starter kit on GitHub, resources for beginners, and organised a Q&A session along with a writing tutorial. Our participants were based in different parts of the world, as shown in Figure 2, with many coming from underrepresented regions. The task consisted of two phases: (1) the development phase and (2) the evaluation phase. During the development phase, the leaderboard was open, allowing a maximum of 999 submissions per participant. In the evaluation phase, the leaderboard was closed, and each participant was allowed up to three submissions, with the last submission being considered for the official ranking.
4.3 Evaluation Metrics and Baselines
Evaluation Metrics
For Tracks A and C, we use the average macro F-score calculated based on the predicted and the gold-standard labels. For Track B, we use the Pearson correlation coefficient, which captures how well the system-predicted intensity scores of test instances align with human judgments. We provided the participants with an evaluation script on our GitHub page.
Our Baselines
We run a simple majority class baseline for each language across all three tracks. Further, for Tracks A and B (Tables 2 and 3, respectively), we fine-tuned RoBERTa using the training data for each language. Table 2 shows the average macro F-scores of the top-performing systems compared to our baseline in Track A, and Table 3 shows the Pearson correlation scores for Track B. For Track C (Table 4), we fine-tuned RoBERTa by training on all languages within a language family while holding out one target language used for testing, e.g., all Indo-European languages except eng when testing on it. For language families with only one language, we trained on the Slavic languages (rus and ukr) and tested on tat; on the Niger-Congo languages (swa and yor) and tested on pcm; and trained on rus when testing on chn.
5 Participating Systems and Results
Lang Team Score Lang Team Score Lang Team Score Lang Team Score
afr pai 0.699 amh chinchunmei 0.773 arq pai 0.669 ary pai 0.629
maomao 0.687 nust titans 0.714 jnlp 0.641 jnlp 0.609
$R_{\text{baseline}}$ 0.371 $R_{\text{baseline}}$ 0.638 $R_{\text{baseline}}$ 0.414 $R_{\text{baseline}}$ 0.472
$M_{\text{baseline}}$ 0.257 $M_{\text{baseline}}$ 0.295 $M_{\text{baseline}}$ 0.445 $M_{\text{baseline}}$ 0.247
chn pai 0.709 deu pai 0.740 eng pai 0.823 esp pai 0.849
teleai 0.682 heimerdinger 0.706 nycu-nlp 0.823 heimerdinger 0.838
$R_{\text{baseline}}$ 0.531 $R_{\text{baseline}}$ 0.642 $R_{\text{baseline}}$ 0.708 $R_{\text{baseline}}$ 0.774
$M_{\text{baseline}}$ 0.278 $M_{\text{baseline}}$ 0.449 $M_{\text{baseline}}$ 0.367 $M_{\text{baseline}}$ 0.312
hau pai 0.751 hin jnlp 0.926 ibo pai 0.600 kin pai 0.657
empaths 0.695 pai 0.920 late-gil-nlp 0.563 mcgill-nlp 0.590
$R_{\text{baseline}}$ 0.596 $R_{\text{baseline}}$ 0.855 $R_{\text{baseline}}$ 0.479 $R_{\text{baseline}}$ 0.463
$M_{\text{baseline}}$ 0.312 $M_{\text{baseline}}$ 0.246 $M_{\text{baseline}}$ 0.236 $M_{\text{baseline}}$ 0.218
mar pai 0.884 orm tewodros 0.616 pcm pai 0.674 ptbr pai 0.683
indidataminer 0.883 late-gil-nlp 0.592 jnlp 0.634 heimerdinger 0.625
$R_{\text{baseline}}$ 0.822 $R_{\text{baseline}}$ 0.126 $R_{\text{baseline}}$ 0.555 $R_{\text{baseline}}$ 0.426
$M_{\text{baseline}}$ 0.264 $M_{\text{baseline}}$ 0.232 $M_{\text{baseline}}$ 0.357 $M_{\text{baseline}}$ 0.2425
ptmz pai 0.548 ron pai 0.794 rus heimerdinger 0.901 som pai 0.577
heimerdinger 0.507 jnlp 0.779 jnlp 0.891 empaths 0.508
$R_{\text{baseline}}$ 0.459 $R_{\text{baseline}}$ 0.762 $R_{\text{baseline}}$ 0.838 $R_{\text{baseline}}$ 0.459
$M_{\text{baseline}}$ 0.163 $M_{\text{baseline}}$ 0.461 $M_{\text{baseline}}$ 0.262 $M_{\text{baseline}}$ 0.198
sun lazarus nlp 0.550 swa empaths 0.386 swe pai 0.626 tat pai 0.846
pai 0.541 pai 0.385 jnlp 0.619 tue-jms 0.797
$R_{\text{baseline}}$ 0.373 $R_{\text{baseline}}$ 0.227 $R_{\text{baseline}}$ 0.520 $R_{\text{baseline}}$ 0.539
$M_{\text{baseline}}$ 0.334 $M_{\text{baseline}}$ 0.179 $M_{\text{baseline}}$ 0.264 $M_{\text{baseline}}$ 0.246
Table 2: Average macro-F1 scores for our baselines ($M_{baseline}$ and $R_{baseline}$ referring to Majority Vote and RoBERTa baselines, respectively) and the top two performing systems for each language in Track A (in bold).
5.1 Overview
Our task attracted more than 700 registered participants and was featured in the Codabench newsletter as the most popular competition hosted on Codabench in 2024.
In the development phase, 153 submissions were made for Track A, 52 for Track B, and 25 for Track C. In the test phase, 220 submissions were made for Track A, 96 for Track B, and 46 for Track C. The official results include more than 220 final submissions from 93 teams. While the English subtracks received the highest number of submissions, we note that other languages, including underserved ones, were comparable in terms of popularity.
We report results only for teams that submitted a system description paper. LABEL:tab:track_a presents the results for Track A, which had 87 participating teams. Table 6 shows the results for Track B, with 38 participating teams, while LABEL:tab:track_c reports the results for Track C, which had 21 participating teams.
5.2 Track A: Multi-label Emotion Detection
5.2.1 Best-Performing Systems
Team Pai
proposes one of the most effective models in the competition. They consistently rank as the top approach in Track A for 20 out of 28 languages. For their system, they combine several base models (ChatGPT-4o, DeepSeek-V3, Gemma-9b, Qwen-2.5-32b, Mistral-Small-24B) using multiple ensemble techniques (neural networks, XGBoost, LightGBM, linear regression, weighted voting). They fine-tune the Gemma and Qwen models using AdaLoRA. For prompting the LLMs, they used an iterative prompt-optimisation technique that generates prompt variations.
Team Chinchunmei
ranks in the top 10 in 16 languages in Track A and 12th in English. They use sample contrastive learning, where performance is enhanced by comparing sample pairs, and generative contrastive learning, where the models learn to distinguish correct from incorrect predictions. Their samples are randomly selected from the task dataset (no external augmentation). They use LLaMa3-Instruct-8B for their fine-tuning.
5.2.2 Takeaways
Most of the teams that rank well on Track A experiment with essentially two methodologies: 1) fine-tuning BERT-based models such as DeBERTa, mBERT, and XLM-R; and/or 2) instruction-fine-tuning using LoRa methodologies in combination with prompt design and data augmentation techniques on LLMs (ChatGPT-4o, DeepSeek-V3, Gemma-9b, Qwen-2.5-32b, Mistral-Small-24B). For instance, Team Telai (3rd in Chinese), Team Empaths (2nd in Hausa and Somali, and 1st in Swahili), Team NYCU-NLP (2nd in English), Team JNLP (1st in Hindi and among the best four across 10 languages), Team Unibouc (1st in Emakhuwa), Team Heiderdinger (2nd in Mozambican Portuguese, German, Spanish, Brazilian Portuguese, and Yoruba; 1st in Russian), and Team Maomao (2nd in Afrikaans).
Few teams focus on only a subset of languages or explore language-related knowledge in their methodology. For instance, Team Lazarus NLP redefined and reformulated the multilabel classification into multiple binary tasks to expand training samples. They also explored how knowledge could potentially be transferred between Indonesian languages. All the top 10 teams performed much better than our baseline model, with an improvement that is more notable in a few low-resource languages, such as Oromo, where Team Tewodros obtained an average macro-F1 score of 0.616 compared to a baseline of 0.126. The same was observed for Yoruba, where Team Pai scored 0.461 compared to a baseline score of 0.092.
5.3 Track B: Emotion Intensity Detection
Lang Team Score Lang Team Score
amh csecu-learners 0.856 arq pai 0.650
heimerdinger 0.781 jnlp 0.587
$R_{\text{baseline}}$ 0.508 $R_{\text{baseline}}$ 0.016
$M_{\text{baseline}}$0.001 $M_{\text{baseline}}$0.009
chn pai 0.722 deu pai 0.766
teleai 0.708 teleai 0.743
$R_{\text{baseline}}$ 0.405 0.562
$M_{\text{baseline}}$ 0.000 $M_{\text{baseline}}$ 0.016
eng pai 0.840 esp pai 0.808
nycu-nlp 0.837 deepwave 0.792
$R_{\text{baseline}}$ 0.641 $R_{\text{baseline}}$ 0.726
$M_{\text{baseline}}$ 0.001 $M_{\text{baseline}}$ 0.011
hau pai 0.770 ptbr pai 0.710
deepwave 0.747 teleai 0.690
$R_{\text{baseline}}$ 0.270 $R_{\text{baseline}}$ 0.297
$M_{\text{baseline}}$ 0.003 $M_{\text{baseline}}$ 0.016
ron pai 0.726 rus pai 0.925
deepwave 0.716 teleai 0.919
$R_{\text{baseline}}$ 0.557 $R_{\text{baseline}}$ 0.877
$M_{\text{baseline}}$ 0.003 $M_{\text{baseline}}$ 0.016
ukr pai 0.708
jnlp 0.672
$R_{\text{baseline}}$ 0.399
$M_{\text{baseline}}$0.01
Table 3: Pearson correlation scores for our baselines (Majority $M_{baseline}$ and RoBERTa $R_{baseline}$) and the top 2 performing systems for each language in Track B (in bold).
5.3.1 Best-Performing Systems
Team Pai
Similar to Track A, Team Pai ranked at the top across all languages in Track B, except for Amharic. They used an ensemble of LLMs, combining several base models (ChatGPT-4oOpenAI (2024), DeepSeek-V3 DeepSeek-AI et al. (2025), Gemma-9b Team et al. (2024), Qwen-2.5-32b Yang et al. (2024), Mistral-Small-24B) with multiple ensemble techniques (neural networks, XGBoost, LightGBM, linear regression, weighted voting). They fine-tuned the Gemma and Qwen models using AdaLoRA. For prompting the LLMs, they employed an iterative prompt-optimisation technique to generate prompt variations.
Team CSECU-Learners
CSECU-Learners ranked at the top in Amharic by fine-tuning language-specific transformers (XLM-Roberta Conneau et al. (2020) for Amharic) with a classification layer and multi-sample dropout.
5.3.2 Takeaways
Teams Deepwave, Teleai, and JNLP also ranked highly across various languages using similar prompt engineering approaches as in Track A. Additionally, Team NYCU-NLP ranked second in English by aggregating instruction-tuned small language models. The different teams outperformed our RoBERTa baseline, which achieved medium Pearson correlation coefficient scores overall, with some extremely low scores in Algerian Arabic, Hausa, Ukrainian, and even Brazilian Portuguese. Overall, we observe that the teams adopted approaches similar to Track A’s, with minor adjustments in the prompts. Furthermore, we observe that even the best-performing teams achieved a Pearson correlation coefficient score of no more than 0.65 on Algerian Arabic, likely due to the novelty of the dataset.
5.4 Track C: Cross-lingual Emotion Detection
Lang Team Score Lang Team Score Lang Team Score Lang Team Score
afr maomao 0.705 amh deepwave 0.661 arq deepwave 0.588 ary deepwave 0.632
deepwave 0.574 uob-nlp 0.627 maomao 0.584 maomao 0.565
$R_{\text{baseline}}$ 0.350 $R_{\text{baseline}}$ 0.487 $R_{\text{baseline}}$ 0.338 $R_{\text{baseline}}$ 0.355
$M_{\text{baseline}}$ 0.257 $M_{\text{baseline}}$ 0.295 $M_{\text{baseline}}$ 0.445 $M_{\text{baseline}}$ 0.247
chn deepwave 0.689 deu deepwave 0.727 eng deepwave 0.797 esp deepwave 0.831
maomao 0.622 gt-nlp 0.687 maomao 0.755 maomao 0.806
$R_{\text{baseline}}$ 0.246 $R_{\text{baseline}}$ 0.468 $R_{\text{baseline}}$ 0.375 $R_{\text{baseline}}$ 0.574
$M_{\text{baseline}}$ 0.278 $M_{\text{baseline}}$ 0.319 $M_{\text{baseline}}$ 0.449 $M_{\text{baseline}}$ 0.367
hau deepwave 0.709 hin deepwave 0.919 ibo deepwave 0.605 ind maomao 0.672
uob-nlp 0.627 maomao 0.896 uob-nlp 0.484 lazarus nlp 0.641
$R_{\text{baseline}}$ 0.320 $R_{\text{baseline}}$ 0.138 $R_{\text{baseline}}$ 0.075 $R_{\text{baseline}}$ 0.376
$M_{\text{baseline}}$ 0.312 $M_{\text{baseline}}$ 0.264 $M_{\text{baseline}}$ 0.236 $M_{\text{baseline}}$
jav heimerdinger 0.439 kin deepwave 0.508 mar deepwave 0.903 orm deepwave 0.542
lazarus nlp 0.438 uob-nlp 0.466 maomao 0.863 uob-nlp 0.491
$R_{\text{baseline}}$ 0.464 $R_{\text{baseline}}$ 0.184 $R_{\text{baseline}}$ 0.772 $R_{\text{baseline}}$ 0.262
$M_{\text{baseline}}$ $M_{\text{baseline}}$ 0.218 $M_{\text{baseline}}$ 0.264 $M_{\text{baseline}}$ 0.232
pcm deepwave 0.674 ptbr deepwave 0.629 ptmz deepwave 0.555 ron deepwave 0.767
maomao 0.562 maomao 0.617 maomao 0.495 maomao 0.747
$R_{\text{baseline}}$ 0.010 $R_{\text{baseline}}$ 0.418 $R_{\text{baseline}}$ 0.297 $R_{\text{baseline}}$ 0.762
$M_{\text{baseline}}$ 0.357 $M_{\text{baseline}}$ 0.243 $M_{\text{baseline}}$ 0.163 $M_{\text{baseline}}$
rus deepwave 0.906 som maomao 0.488 sun deepwave 0.467 swa maomao 0.381
maomao 0.852 deepwave 0.488 maomao 0.464 deepwave 0.355
$R_{\text{baseline}}$ 0.704 $R_{\text{baseline}}$ 0.273 $R_{\text{baseline}}$ 0.194 $R_{\text{baseline}}$ 0.190
$M_{\text{baseline}}$ 0.262 $M_{\text{baseline}}$ 0.198 $M_{\text{baseline}}$ 0.334 $M_{\text{baseline}}$ 0.179
swe deepwave 0.645 tat deepwave 0.789 tir deepwave 0.505 ukr deepwave 0.702
maomao 0.578 maomao 0.697 uob-nlp 0.445 maomao 0.623
$R_{\text{baseline}}$ 0.512 $R_{\text{baseline}}$ 0.445 $R_{\text{baseline}}$ 0.339 $R_{\text{baseline}}$ 0.496
$M_{\text{baseline}}$ 0.264 $M_{\text{baseline}}$ 0.246 $M_{\text{baseline}}$ 0.253 $M_{\text{baseline}}$ 0.157
vmw deepwave 0.210 xho maomao 0.443 yor maomao 0.360 zul maomao 0.397
ozemi 0.193 ozemi 0.315 deepwave 0.342 heimerdinger 0.226
$R_{\text{baseline}}$ 0.052 $R_{\text{baseline}}$ 0.127 $R_{\text{baseline}}$ 0.053 $R_{\text{baseline}}$ 0.153
$M_{\text{baseline}}$ 0.162 $M_{\text{baseline}}$ $M_{\text{baseline}}$ 0.165 $M_{\text{baseline}}$
Table 4: Average macro-F1 scores for our baselines ($M_{baseline}$ and $R_{baseline}$ referring to Majority Vote and RoBERTa baselines, respectively) and the top 2 performing systems for each language in Track C (in bold).
5.4.1 Best-Performing Systems
Team deepwave
Team Deepwave fine-tuned Google Gemma-2 using tailored data augmentation and Chain-of-Thought (CoT) prompting. They decomposed the task into two sub-tasks: (1) sentiment keyword identification and (2) sentiment polarity recognition.
To address the challenge of limited data, they employed k-fold (k=5) cross-validation and used model merging—a strategy that combines the predictions of multiple models to improve generalization—by averaging the prediction probabilities of each model, assigning equal weights of 0.2. In this track, a dedicated LoRA module was trained for each target language. The training dataset for each module comprised data from all other languages in Track A, excluding the target language $L_{i}$. They exclusively used augmented data generated through CoT prompting for training.
Team maomao
Team maomao experimented with different setups for fine-tuning LLMs. The base models used were Qwen2.5-7B-Instruct, GPT-0544o Mini2, and LLaMA-3.2-3B-Instruct (Touvron et al., 2023). They applied Direct Preference Optimization to refine their model—a technique that selects high-quality instances from lower-quality ones within a dataset. After this step, they retrained the model using the refined dataset. They also explored random sampling and retrieval-augmented generation (RAG) methods for training, primarily on DeepSeek-V3, Qwen-Max5093, and Grok-V2.
Figure 3: Top LLMs used by participants across tracks (A, B, and C).
5.4.2 Takeaways
Other top-performing systems include Team Ozemi, who fine-tuned a multilingual BERT model and applied machine translation to enhance performance across 35 languages. They used the Synthetic Minority Oversampling Technique (SMOTE) with TF-IDF to address class imbalance in Russian. It appears they translated all datasets into a common language using Google Translate before processing, except for Nigerian Pidgin and Emakhua, where they used a multilingual BERT model for translation. They also leveraged two Kaggle competition datasets for data augmentation. Team heimerdinger, who ranked highest in Javanese, built their approach using various LLMs (LLaMA 3.1 8B, Qwen 2.5 7B, DeepSeek-7B, MistralV0.3-7B, and Gemma2-9B) for Track C. They employed in-context learning with multilingual examples from high-resource languages such as English and Spanish.
Overall, the different teams outperformed our baselines. However, we observe that the average scores for this track are lower, particularly in low-resource languages such as Emakhuwa, as participants were presented with an additional challenge.
6 Discussion
Popular Methods
Unsurprisingly, most top-performing teams favored fine-tuning and prompting large language models (LLMs) such as Gemma-2, Mistral, Phi-4, Qwen-2.5, DeepSeek, LLaMA-3, GPT, and Gemini models. For fine-tuning, both full fine-tuning and parameter-efficient fine-tuning were the most commonly used strategies to enhance performance.
For prompting, few-shot, zero-shot, and chain-of-thought prompting were the most frequently used techniques.
Many participants also experimented with traditional transformer-based models, particularly XLM-RoBERTa, mBERT, DeBERTa, and IndicBERT (see Figure 3 and Figure 4 in the Appendix).
Best Performing Systems
The results from the top-performing submissions suggest that while LLMs achieve strong overall performance, their effectiveness heavily depends on prompt engineering techniques and wording.
Additionally, performance varies significantly by language. Across all tracks, LLM-based approaches and the best-performing systems consistently yielded better results for high-resource languages such as English and Russian. Conversely, performance dropped notably when tested on low-resource languages such as Swahili and Emakhuwa.
Furthermore, most teams did not incorporate additional datasets to enhance performance (see Appendix) as few-shot and zero-shot approaches proved highly effective.
7 Conclusion
We presented our shared task on text-based emotion recognition, which covered three tracks and a total of 32 languages. The submitted systems were ranked based on macro F1-scores for Tracks A and C, comparing predicted labels to gold labels, and based on the ranking of predicted intensity scores for Track B. We summarised the reported results, highlighting the best-performing and most innovative methods. Overall, performance varied significantly across languages.
Limitations
Emotions are subjective and subtle, and they are expressed and perceived differently. We do not claim that our datasets capture the true emotions of the speakers, fully represent language use across the 32 languages, or cover all possible emotions. We discuss these limitations extensively in Ethical Considerations.
We acknowledge the limited data sources available for some low-resource languages. Therefore, our datasets may not be suitable for tasks requiring large amounts of data in a given language. However, they serve as a valuable starting point for research in this area.
Ethical Considerations
Emotion perception and expression are subjective and nuanced, as they are influenced by various factors (e.g., cultural background, social group, personal experiences, and social context). Thus, it is impossible to determine someone’s emotions with absolute certainty based solely on short text snippets. Our datasets explicitly focus on perceived emotions—identifying the emotions that most people believe the speaker may have felt. We do not claim to annotate the speaker’s true emotions, as these cannot be definitively determined from text alone. We recognise the importance of this distinction, as perceived emotions may differ from actual emotions.
We acknowledge potential biases in our data, as we rely on text-based communication, which inherently carries biases from data sources and annotators. Additionally, while many of our datasets focus on low-resource languages, we do not claim they fully represent these languages’ usage. Further, although we took measures to filter inappropriate content, some instances may have been overlooked.
We explicitly urge careful consideration of ethical implications before using our datasets. We prohibit their use for commercial purposes or by state actors in high-risk applications unless explicitly approved by the dataset creators. Systems built on our datasets may not be reliable at the individual instance level and are susceptible to domain shifts. Thus, they should not be used for critical decision-making, such as in health applications, without expert supervision. For a more in-depth discussion, see Mohammad (2022, 2023).
Finally, all annotators involved in the study were compensated above the minimum hourly wage.
Acknowledgments
Shamsuddeen Muhammad acknowledges the support of Google DeepMind and Lacuna Fund, an initiative co-founded by The Rockefeller Foundation, Google.org, and Canada’s International Development Research Centre. The views expressed herein do not necessarily represent those of Lacuna Fund, its Steering Committee, its funders, or Meridian Institute.
Nedjma Ousidhoum would like to thank Abderrahmane Samir Lazouni, Lyes Taher Khalfi, Manel Amarouche, Narimane Zahra Boumezrag, Noufel Bouslama, Abderaouf Ousidhoum, Sarah Arab, Wassil Adel Merzouk, Yanis Rabai, and another annotator who would like to stay anonymous for their work and insightful comments.
Idris Abdulmumin gratefully acknowledges the ABSA UP Chair of Data Science for funding his post-doctoral research and providing compute resources.
Jan Philip Wahle and Terry Ruas were partially supported by the Lower Saxony Ministry of Science and Culture and the VW Foundation.
Meriem Beloucif acknowledges Nationella Språkbanken (the Swedish National Language Bank) and Swe-CLARIN – jointly funded by the Swedish Research Council (2018–2024; dnr 2017-00626) and its 10 partner institutions for funding the Swedish annotations.
Alexander Panchenko would like to thank Nikolay Ivanov, Artem Vazhentsev, Mikhail Salnikov, Maria Marina, Vitaliy Protasov, Sergey Pletenev, Daniil Moskovskiy, Vasiliy Konovalov, Elisey Rykov, and Dmitry Iarosh for their help with the annotation for Russian. Preparation of Tatar data was funded by AIRI and completed by Dina Abdullina, Marat Shaehov, and Ilseyar Alimova.
Yi Zhou would like to thank Gaifan Zhang, Bing Xiao, and Rui Qin for their help with the annotations and for providing feedback.
Appendix A Task Results and Analysis
In this section, we provide the complete shared task results in LABEL:tab:track_a, LABEL:, 6, LABEL: and LABEL:tab:track_c, and some statistics on the architectures, training strategies, fine-tuning and prompting techniques used by our participants ( Figure 4). We also provide information on additional training data usage Figure 5.
Table 5: Track A Results
S/N Team Name afr amh arq ary chn deu eng esp hau hin ibo kin mar orm pcm ptbr ptmz ron rus som sun swa swe tat tir ukr vmw yor avg
1 PAI 69.9 64.7 66.9 62.9 70.9 74.0 82.3 84.9 75.1 92.0 60.0 65.7 88.4 58.2 67.4 68.3 54.8 79.4 88.2 57.6 54.1 38.5 62.6 84.6 53.8 72.6 25.5 46.1 66.8
2 University of Indonesia 54.6 51.2 55.0 53.5 62.9 66.2 74.9 79.5 64.4 86.1 49.6 48.4 84.7 46.2 60.5 56.9 42.7 73.8 84.4 43.7 43.2 37.9 57.8 60.4 40.0 63.4 13.6 29.0 56.6
3 LATE-GIL-NLP 35.0 58.1 48.6 51.3 57.4 63.3 76.3 75.2 61.5 87.3 56.3 51.1 78.8 59.2 51.7 54.7 47.5 70.8 79.8 36.9 40.8 28.5 46.0 65.5 58.7 47.5 16.3 37.5 55.1
4 Zero_Shot 27.0 54.1 23.9 23.1 46.7 46.7 69.5 73.6 55.8 83.8 47.0 19.8 80.7 28.6 42.9 31.3 12.6 64.5 84.3 30.9 26.5 15.8 44.5 40.6 31.3 44.6 0.0 12.5 41.5
5 NTA 40.7 67.2 50.6 24.9 59.4 57.1 76.1 75.3 67.8 83.7 47.0 38.0 81.4 50.5 47.8 34.1 34.6 70.0 83.5 47.1 42.0 21.4 52.9 56.9 59.1 53.3 20.8 21.9 52.3
6 XLM-Muriel 42.7 66.4 46.1 50.9 59.3 58.4 62.8 74.6 58.8 84.0 47.8 32.1 84.2 50.8 53.1 51.6 43.2 65.9 84.2 44.6 44.1 27.1 58.6 67.5 48.2 61.9 16.2 30.2 54.1
7 PromotionGo 22.1 42.7 40.7 37.9 39.4 43.2 65.5 60.1 53.4 66.4 47.6 31.3 64.5 45.8 38.6 32.8 30.9 61.7 62.7 35.2 32.1 15.1 36.8 52.0 38.3 31.8 13.0 21.1 41.5
8 OZemi 46.9 37.0 48.0 34.0 49.9 50.8 63.8 52.5 37.6 46.8 28.5 29.9 49.8 31.1 46.4 34.6 24.2 64.2 70.6 30.9 39.9 23.4 38.3 40.6 33.1 27.9 19.3 21.1 40.0
9 INFOTEC-NLP 50.9 69.4 51.7 55.5 64.8 66.3 70.9 79.2 64.5 89.4 50.0 35.3 88.0 55.1 57.9 55.9 48.5 74.0 87.2 48.9 42.9 28.1 59.4 68.1 49.3 66.0 4.0 15.3 57.0
10 Pixel Phantoms 30.6 43.7 32.1 34.2 39.6 37.9 59.7 64.0 50.1 73.0 41.0 31.7 70.2 31.9 40.9 27.5 20.8 57.6 72.0 25.4 30.9 17.8 38.9 40.6 32.0 35.3 7.8 14.7 39.4
11 YNWA_PZ 54.0 61.2 51.1 51.9 56.6 60.6 74.0 76.2 63.2 80.3 50.9 51.9 81.1 54.3 53.1 48.0 50.1 73.8 82.4 48.3 42.5 29.5 56.5 64.3 52.4 48.6 16.8 34.1 56.0
12 Deepwave 52.3 50.7 48.2 52.4 62.0 65.5 75.1 79.4 59.4 90.1 48.1 38.0 87.8 44.6 59.5 53.9 45.1 70.3 87.2 30.9 37.9 22.7 56.9 68.2 40.9 61.9 7.8 24.0 54.3
13 UB_Tel-U 55.1 48.4 52.9 43.3 50.6 58.3 70.3 70.8 51.6 66.7 38.6 32.9 68.4 35.9 53.8 47.1 36.7 69.2 65.5 29.9 41.9 30.2 50.6 54.6 40.5 42.1 14.0 22.5 47.9
14 Heimerdinger 57.4 55.8 59.5 58.2 68.0 70.6 80.4 83.8 61.5 89.6 54.1 44.5 87.7 50.9 60.5 62.5 50.7 74.6 90.1 43.8 48.2 30.2 58.1 75.4 45.9 63.7 22.2 39.2 60.3
15 Chinchunmei 57.7 77.3 57.2 56.3 66.8 70.4 79.7 81.7 67.6 90.0 55.9 47.0 87.2 53.7 63.3 60.3 50.3 74.7 88.7 43.4 45.4 31.9 58.5 75.8 48.0 65.8 19.7 28.9 60.8
16 TeleAI 57.8 37.6 39.8 52.8 68.2 67.2 80.6 80.7 53.8 85.2 38.5 39.5 78.3 38.1 54.6 59.9 44.9 68.2 87.7 35.2 46.6 32.0 57.3 52.9 37.0 62.3 9.0 32.2 53.5
17 FiRC-NLP 58.0 69.6 55.2 58.5 61.7 68.4 74.2 79.7 69.4 87.8 54.1 41.5 86.7 60.7 61.1 60.1 51.1 74.9 86.6 55.9 49.8 32.3 60.9 71.7 57.6 67.3 25.0 30.8 61.1
18 maomao 68.7 49.1 45.4 54.8 60.3 57.0 63.2 77.8 55.7 81.9 32.0 43.1 83.7 28.4 54.7 61.2 46.4 71.1 83.1 42.3 45.3 31.3 56.0 72.8 38.0 63.3 16.7 34.3 54.2
19 McGill-NLP 60.1 63.4 58.7 52.5 59.6 64.4 77.7 79.0 65.4 88.0 53.2 59.0 87.5 56.4 63.1 56.8 45.1 72.8 86.4 48.4 50.3 35.9 51.1 73.9 48.2 60.0 12.6 33.0 59.4
20 Empaths 66.7 55.8 52.0 37.0 62.5 21.6 78.8 81.7 69.5 86.6 53.8 56.3 86.4 49.0 59.9 37.2 49.3 76.1 87.9 50.8 52.9 38.6 49.3 77.6 20.4 61.6 21.6 38.2 56.4
21 Pateam 26.6 33.3 56.3 44.5 58.0 43.8 80.5 82.8 36.7 78.2 26.0 21.9 67.5 40.6 33.7 42.7 49.2 41.3 61.3 21.1 42.3 - 37.6 39.7 10.5 50.4 6.7 13.9 42.5
22 UoB-NLP 45.0 67.0 46.9 51.5 56.9 54.8 64.5 72.9 66.3 84.2 47.2 51.6 80.0 52.6 51.5 47.8 40.0 68.1 - 44.5 36.5 26.2 52.1 63.7 50.3 58.4 5.6 19.3 52.1
23 YNU-NPCC 43.5 47.6 44.4 48.4 55.8 60.3 - 73.1 50.0 82.2 36.6 44.3 82.0 40.1 44.5 53.2 38.6 60.2 79.9 38.2 49.0 31.3 52.2 70.5 28.2 56.9 4.6 26.1 49.7
24 JNLP 59.2 67.7 64.1 60.9 68.0 - 80.4 83.0 65.0 92.6 54.0 42.9 87.8 57.3 63.4 61.8 45.4 77.9 89.1 49.6 46.0 29.5 61.9 72.2 48.5 - 22.6 36.1 61.0
25 CSIRO-LT 31.6 58.8 52.5 - 57.8 65.7 77.6 82.0 54.2 73.2 32.6 43.0 75.8 39.2 21.8 55.5 - 72.2 89.1 43.3 48.8 30.3 48.9 53.8 50.0 66.4 17.3 30.0 52.7
26 Howard University-AI4PC - 15.4 30.4 32.3 39.3 40.1 - 57.0 31.8 62.9 18.5 16.6 58.4 17.5 29.5 38.6 28.2 43.8 59.0 16.0 32.3 22.9 30.7 30.1 - 35.3 12.8 13.6 32.5
27 Tue-JMS - 69.6 53.4 58.0 60.3 64.5 68.5 79.4 69.0 88.5 53.0 53.2 87.3 50.9 56.9 56.5 47.8 73.7 88.0 47.8 43.9 - 59.0 79.7 53.3 63.4 - - 63.6
28 SyntaxMind 36.5 - 45.7 37.3 55.8 48.7 66.5 57.4 - 65.1 - - 72.5 - - 31.4 37.1 61.7 66.0 - 35.6 24.1 43.3 49.1 - 31.5 - 26.1 46.9
29 IASBS 59.1 52.3 - - 55.4 61.4 73.5 72.8 65.8 81.1 54.4 - 80.1 - - - - 72.5 82.2 - - - 53.0 - - 58.8 - - 65.9
30 Emotion Train - - 56.5 56.8 - 65.2 75.6 76.1 - - - - - - 59.3 53.6 47.1 72.2 88.2 - - 12.4 53.2 49.8 - 45.3 - - 58.0
31 Team A - 68.5 - 55.5 61.2 66.5 73.4 80.5 67.7 89.0 - - 86.6 56.3 - - - 74.0 88.3 - - - - - - 60.1 - - 71.4
32 Habib University 25.7 45.5 - 14.5 - 44.6 71.0 71.3 - 80.8 - - 76.4 - - - - 50.4 72.1 - - - 21.9 - - 29.1 - - 50.3
33 JellyK - - - - 63.9 65.4 76.2 77.2 - 86.1 - - 86.3 - 58.2 - 42.2 74.2 88.9 - - - - 59.3 - 57.9 - - 69.6
34 CSECU-Learners - 70.2 55.5 - 60.0 60.2 73.8 76.9 67.3 - - - - - - 52.4 - 74.7 84.7 - - - - - - 50.6 - - 66.0
35 UncleLM - - - 53.9 61.5 58.2 74.6 76.4 31.8 85.6 - - 80.0 - - - 44.0 74.5 87.1 - - - - - - - - - 66.1
36 Domain_adaptation - 47.8 26.4 - 57.2 57.7 - 72.0 30.2 - - - - - - 49.3 - 54.0 76.0 - - - - - - 50.4 - - 52.1
37 Team KiAmSo 46.7 53.4 49.1 50.1 56.6 59.8 72.5 76.3 58.4 78.5 - - - - - - - - - - - - - - - - - - 60.1
38 EMO-NLP - 60.2 - - - 53.1 66.8 - - - - - - 31.8 - 45.9 - - 75.9 37.8 29.3 - - - 32.5 - - - 48.1
39 GT-NLP 51.9 - 52.7 53.1 - 66.3 - - - - - - - - - 59.9 - - 86.4 - - - 56.4 - - 60.0 - - 60.8
40 Trans-Sent - 54.7 - - - 48.0 70.4 - - 82.4 - - 80.3 - - - - 70.9 88.3 - - - - - - - - - 70.7
41 NUST Titans - 71.4 - - - - - - 69.2 - 18.1 - - - - - - - - - 34.5 29.5 - - - - - 31.2 42.3
42 Zero - - - - - - 71.7 35.9 - - - - - - - - - 36.7 52.9 - - - - - - - - - 49.3
43 Tewodros - 70.0 - - - - 71.2 79.4 - - - - - 61.6 - - - - - - - - - - - - - - 70.6
44 UIMP-Aaman - - - - - 62.4 72.2 76.7 - - - - - - - 51.7 - - - - - - - - - - - - 65.8
45 GinGer 45.4 - 54.7 - - - - - - 83.9 - - - - - - - - - - - - 52.3 - - - - - 59.1
46 DUT_IR - - - - - - 81.2 - - 84.2 - - 83.7 - - - - - 74.4 - - - - - - - - - 80.9
47 DataBees - - - - - 51.7 37.2 69.6 - - - - - - - - - - - - - - - - - - - - 52.8
48 Team Unibuc - - - - - - 75.5 - - - - - - - - - 17.3 - - - - - - - - - 32.5 - 41.8
49 indiDataMiner - - - - - - 66.3 - - 87.4 - - 88.3 - - - - - - - - - - - - - - - 80.7
50 CIOL - - - - - - 62.1 - - - - - - - - 27.7 - - 84.8 - - - - - - - - - 58.2
51 CIC-IPN - - - - - - 70.6 73.2 - - - - - - - - - - - - - - - - - - - 28.2 57.3
52 Team UBD - - - - - 54.3 68.0 76.2 - - - - - - - - - - - - - - - - - - - - 66.2
53 ITF-NLP - - - - - 62.2 75.0 - - - - - - - - - - - - - - - - - - - - - 68.6
54 IEGPS-CSIC - - - - - - 71.0 - - - - - - - - - - - 84.2 - - - - - - - - - 77.6
55 NLP-DU - - - - - - 70.6 - - - - - - - - - - - - - - - - - - - - - 70.6
56 zhouyijiang1 - - - - - - 70.5 - - - - - - - - - - - - - - - - - - - - - 70.5
57 wangkongqiang - - - - - - 70.0 - - - - - - - - - - - - - - - - - - - - - 70.0
58 JUNLP - - - - - - 69.9 - - - - - - - - - - - - - - - - - - - - - 69.9
59 dianchi - - - - - - 69.6 - - - - - - - - - - - - - - - - - - - - - 69.6
60 RSSN - - - - - - 69.4 - - - - - - - - - - - - - - - - - - - - - 69.4
61 CodeBlAIzers - - - - - - 55.2 - - - - - - - - - - - - - - - - - - - - - 55.2
62 Angeliki Linardatou - - - - - - 67.8 - - - - - - - - - - - - - - - - - - - - - 67.8
63 TILeN - - - - - - - - - - - - - - - - - - 82.1 - - - - - - - - - 82.1
64 AKCIT - - - - - - - - - - - - - - - 60.2 - - - - - - - - - - - - 60.2
65 STFXNLP - - - - - - 64.7 - - - - - - - - - - - - - - - - - - - - - 64.7
66 VerbaNexAI - - - - - - 62.7 - - - - - - - - - - - - - - - - - - - - - 62.7
67 NLP-Cimat - - - - - - 62.4 - - - - - - - - - - - - - - - - - - - - - 62.4
68 HausaNLP - - - - - - - - 56.0 - - - - - - - - - - - - - - - - - - - 56.0
69 TechSSN3 - - - - - - 69.0 - - - - - - - - - - - - - - - - - - - - - 69.0
70 HTU - - 48.6 - - - - - - - - - - - - - - - - - - - - - - - - - 48.6
71 xiacui - - - - - - 72.2 - - - - - - - - - - - - - - - - - - - - - 72.2
72 GOLDX - - - - - - 76.2 - - - - - - - - - - - - - - - - - - - - - 76.2
73 Tarbiatmodares - 65.5 - - - - - - - - - - - - - - - - - - - - - - - - - - 65.5
74 Amado - 63.0 - - - - - - - - - - - - - - - - - - - - - - - - - - 63.0
75 AtlasIA - - - 33.5 - - - - - - - - - - - - - - - - - - - - - - - - 33.5
76 NYCU-NLP - - - - - - 82.2 - - - - - - - - - - - - - - - - - - - - - 82.2
77 NCL-NLP - - - - - - 78.6 - - - - - - - - - - - - - - - - - - - - - 78.6
78 MRS - - - - - - 76.6 - - - - - - - - - - - - - - - - - - - - - 76.6
79 NITK-VITAL - - - - - - 75.6 - - - - - - - - - - - - - - - - - - - - - 75.6
80 CharsiuRice - - - - - - 72.6 - - - - - - - - - - - - - - - - - - - - - 72.6
81 QiMP - - - - - - 75.4 - - - - - - - - - - - - - - - - - - - - - 75.4
82 NLP_goats - - - - - - 75.0 - - - - - - - - - - - - - - - - - - - - - 75.0
83 AGHNA - - - - - - 73.7 - - - - - - - - - - - - - - - - - - - - - 73.7
84 OPI-DRO-HEL at - - - - - - 73.6 - - - - - - - - - - - - - - - - - - - - - 73.6
85 Exploration Lab IITK - - - - - - 73.4 - - - - - - - - - - - - - - - - - - - - - 73.4
86 Lotus - - - - - - 73.2 - - - - - - - - - - - - - - - - - - - - - 73.2
87 Lazarus NLP - - - - - - - - - - - - - - - - - - - - 55.0 - - - - - - - 55.0
S/N Team Name amh arq chn deu eng esp hau ptbr ron rus ukr avg
1 CSECU-Learners 85.6 44.3 57.1 53.3 65.0 71.5 65.6 46.6 63.7 83.3 47.8 62.2
2 Heimerdinger 78.0 50.7 67.1 72.3 79.3 78.8 61.8 65.1 65.7 88.3 60.3 69.8
3 CSIRO-LT 15.4 52.1 54.9 61.0 72.1 69.7 37.1 48.9 59.2 79.3 49.4 54.5
4 TeleAI 32.2 52.4 70.8 74.2 83.2 78.6 57.9 69.0 70.4 91.8 64.5 67.7
5 Pixel Phantoms 37.7 11.2 36.6 34.2 32.9 52.8 46.7 25.0 40.6 74.5 24.8 37.9
6 IASBS 46.1 49.8 53.6 59.8 69.3 69.2 65.8 52.1 62.0 86.0 56.5 60.9
7 maomao 46.2 56.7 58.6 65.4 75.2 74.9 52.0 60.2 65.3 84.7 53.4 63.0
8 YNWA_PZ 49.4 36.5 48.5 54.1 68.8 66.7 58.4 38.2 57.6 78.4 42.5 54.5
9 UB_Tel-U 56.9 33.2 56.6 55.1 53.1 68.3 53.0 47.8 55.6 78.2 53.6 55.6
10 Habib University 60.3 30.8 47.8 53.2 74.0 64.2 39.5 38.6 49.9 76.4 42.5 52.5
11 JNLP 60.4 58.7 65.9 72.5 81.3 77.5 65.0 65.1 70.5 90.7 67.2 70.4
12 PAI 64.6 65.0 72.2 76.6 84.0 80.8 77.0 71.0 72.6 92.5 70.8 75.2
13 Tue-JMS 67.2 42.5 61.3 64.3 66.5 73.9 67.0 56.0 65.4 88.6 56.1 64.4
14 Deepwave 68.0 57.4 69.4 74.2 80.9 79.2 74.7 68.4 71.6 91.6 66.2 72.9
15 Chinchunmei 71.6 48.6 66.4 70.8 81.0 77.0 67.9 62.2 66.4 89.8 63.1 69.5
16 Domain_adaptation 41.3 20.6 3.4 30.1 - 45.5 54.6 23.9 36.8 56.1 21.1 33.3
17 GT-NLP - 52.8 57.7 49.1 75.1 75.8 63.3 61.6 68.5 88.0 31.9 62.4
18 SyntaxMind - 15.8 47.9 38.9 55.4 39.2 - 23.6 36.8 52.6 19.1 36.6
19 Zero_Shot - 29.2 46.7 42.3 - 63.6 57.3 - 51.8 83.7 - 53.5
20 Stanford MLab - 21.9 - - 69.0 - 54.0 - 39.4 75.9 - 52.0
21 UT-NLP - - 30.7 - 56.9 - 33.7 - - - 35.2 39.1
22 UIMP-Aaman - - - 61.6 73.3 70.5 - 50.4 - - - 64.0
23 Tewodros 65.0 - - - 59.4 62.3 - - - - - 62.2
24 CIOL - - 48.3 - - 68.5 - - - 85.9 - 67.6
25 Team UBD - - - 48.9 - 70.6 - - - - - 59.8
26 Zero - - - - 61.3 - - - - - - 61.3
27 AKCIT - - - - - - - 64.8 - - - 64.8
28 DataBees - - - - 3.0 - - - - - - 3.0
29 QiMP - - - - 49.2 - - - - - - 49.2
30 CharsiuRice - - - - 56.6 - - - - - - 56.6
31 OZemi - - - - 61.2 - - - - - - 61.2
32 tinaal - - - - 67.0 - - - - - - 67.0
33 CIC-IPN - - - - 63.9 - - - - - - 63.9
34 indiDataMiner - - - - 69.0 - - - - - - 69.0
35 NLP_goats - - - - 75.1 - - - - - - 75.1
36 AGHNA - - - - 76.2 - - - - - - 76.2
37 NYCU-NLP - - - - 83.7 - - - - - - 83.7
38 TILeN - - - - - - - - - 80.7 - 80.7
Table 6: Track B Results
Table 7: Track C Results
S/N Team Name afr amh arq ary chn deu eng esp hau hin ibo ind jav kin mar orm pcm ptbr ptmz ron rus som sun swa swe tat tir ukr vmw xho yor zul avg
1 maomao 70.5 58.0 58.4 56.5 62.2 68.3 75.4 80.6 59.1 89.6 43.3 67.2 42.2 44.9 86.3 36.0 56.2 61.7 49.5 74.7 85.2 48.8 46.4 38.0 57.8 69.7 37.9 62.3 17.3 44.3 35.9 39.7 57.0
2 YNWA_PZ 54.0 61.2 51.1 51.9 56.6 60.6 74.0 76.2 63.2 80.3 50.9 35.6 25.6 51.9 81.1 54.3 53.1 48.0 50.1 73.8 82.4 48.3 42.5 29.5 56.5 64.3 52.4 48.6 16.8 16.6 34.1 16.4 51.9
3 Deepwave 57.4 66.1 58.8 63.2 68.9 72.7 79.7 83.1 70.9 91.9 60.5 55.4 37.2 50.8 90.3 54.2 67.4 62.9 55.5 76.7 90.6 48.8 46.7 35.5 64.5 78.9 50.5 70.2 21.0 15.9 34.2 19.3 59.4
4 OZemi 47.1 37.0 47.9 33.9 49.9 50.8 63.8 52.7 37.8 46.7 27.6 48.8 41.3 29.9 49.7 31.2 46.4 34.5 24.2 64.5 44.0 31.0 40.8 23.2 38.8 40.4 33.1 28.1 19.3 31.5 21.1 21.2 38.7
5 UB_Tel-U 37.1 62.3 42.3 44.5 58.8 60.3 65.6 74.8 58.6 85.8 43.0 51.2 35.3 29.1 83.3 37.6 52.8 49.9 37.8 70.3 83.1 35.1 37.5 20.2 54.5 63.9 36.4 57.9 4.2 16.3 13.9 10.8 47.3
6 UoB-NLP 36.3 62.7 44.1 42.4 56.9 55.4 64.5 72.9 62.7 80.0 48.4 33.3 - 46.6 80.0 49.1 51.3 44.0 32.5 68.1 80.6 39.7 29.6 22.0 48.3 63.7 44.5 49.8 3.4 - 19.3 - 49.4
7 CIOL 15.7 - - - 29.9 - 22.3 26.2 - 22.7 - 26.1 24.4 - 17.9 - - - 9.8 23.1 31.0 - 22.1 19.0 21.3 23.0 - 19.7 - - 15.1 - 21.7
8 IASBS 32.8 - 48.5 - 54.3 47.2 73.0 69.6 59.8 76.4 - - - - 77.4 - - 41.0 43.5 64.9 81.7 - - - 47.3 - - 50.9 - - - - 57.9
9 Habib University 30.0 27.0 - - 39.1 8.0 - 24.1 - 67.4 9.2 - - - 76.4 - - - - 47.6 20.7 27.9 - 11.6 22.1 - - 17.9 - - - - 30.6
10 Howard University-AI4PC - 14.0 41.8 29.5 - - - - - 50.1 25.1 - 37.5 20.6 - 17.6 - - 23.3 51.3 - 16.3 - 21.2 - - 17.9 - - - - - 28.2
11 Domain_adaptation - 47.2 40.2 - 6.4 32.2 46.2 56.1 50.1 - - - - - - - - 25.2 - 51.9 54.7 - - - - - - 23.5 - - - - 39.4
12 Zero_Shot - 45.9 43.6 30.3 - - - - 49.9 - - - - - - 39.1 37.5 - 20.2 - - 36.5 29.6 18.4 - - - - - - - - 35.1
13 GT-NLP 53.8 - 54.3 53.5 - 68.7 - - - 78.3 - 58.3 - - - - - - - - 80.6 - - - 57.5 - - 55.6 - - - - 62.3
14 UT-NLP - - - - 33.7 - 56.0 - 39.3 - - 50.5 - - - - - - - - - - - - - - - 23.7 - - - 18.9 37.0
15 UncleLM - - - - - - - 60.9 - 70.9 - - - - 78.3 - - - - - 67.0 - - - - 36.8 - 55.9 - - - - 61.6
16 Team UBD - - - - - - - - - 60.1 - 41.0 - - - - - 40.5 - 50.9 53.4 - - - - - - 38.6 - - - - 47.4
17 Tue-JMS 31.3 - - - - - - - - - - 50.8 34.7 - - - - - - - - - - - - - - - - 10.8 - 13.1 28.1
18 UIMP-Aaman - - - - - 59.4 66.1 69.4 - - - - - - - - - 49.4 - - - - - - - - - - - - - - 61.1
19 Heimerdinger - - - - - - - - - - - 60.9 43.9 - - - - - - - - - - - - - - - - 26.0 - 22.6 38.4
20 Lazarus NLP - - - - - - - - - - - 64.1 43.8 - - - - - - - - - - - - - - - - - - - 54.0
21 GOLDX - - - - - - - - - - - 42.9 - - - - - - - - - - - - - - - - - - - - 42.9
Figure 4: Overview of the most common architectures, training techniques, fine-tuning strategies, and prompting strategies used by our participants.
Figure 5: Overview of additional data and compute resources used by our participants.
Table 8: Participants Information
Tasks Team Affiliation Paper
A, B AGHNA Institut Teknologi Bandung Abyan (2025)
A, B, C AKCIT CEIA NLP Brito et al. (2025)
A Amado none Bade et al. (2025)
A Angeliki Linardatou none Linardatou and Platanou (2025)
A AtlasIA AtlasIA Majjodi et al. (2025)
A, B CIC-NLP none Abiola et al. (2025)
A, B, C CIOL Shahjalal University of Science Technology, Bangladesh Hoque et al. (2025)
A, B CSECU-Learners University of Chittagong Ahmad et al. (2025)
A, B CSIRO-LT CSIRO, Macquarie University Chen et al. (2025a)
A, B CharsiuRice Eberhard Karls University of Tübingen Yip et al. (2025)
A, B Chinchunmei Newcastle University, Shumei AI Research Institute Li et al. (2025)
A CodeBlAIzers Queen Mary University of London Wang and Gupta (2025)
A DUT_IR Dalian University of Technology, China Liu et al. (2025)
A DataBees SSN College Of Engineering Anand et al. (2025)
A, B, C Deep none Shenpo (2025)
A, B Domain_adaptation Moscow Institute of Physics and Technology, University of Leeds Lepekhin and Sharoff (2025)
A EMO-NLP Yunnan University Li and Yang (2025)
A Emotion Train Mohamed bin Zayed University of Artificial Intelligence Demidova et al. (2025)
A Empaths HSE University, Independent researcher Morozov et al. (2025)
A Exploration Lab IITK IIT Kanpur Nadeem et al. (2025)
A FiRC-NLP none Tufa et al. (2025)
A, C GOLDX none Clinton (2025)
A, B, C GT-NLP Georgia Institute of Technology, Johnson & Johnson, Novastraum LLC, C3 AI Saeedi et al. (2025)
A, B GinGer Sharif University of Technology, Iran University of Science and Technology Naebzadeh and Askari (2025)
A HTU AlHussein Technical University Saleh and Biltawi (2025)
A, B, C Habib University Habib University, Karachi, Pakistan Waheed et al. (2025)
A HausaNLP University of the Witwatersrand Johannesburg, Federal Polytechnic, Daura Sani et al. (2025)
A, B, C Heimerdinger Huazhong University of Science and Technology Tong et al. (2025)
A, C Howard University-AI4PC Howard University Ince and Aryal (2025)
A, B, C IASBS IASBS University Tareh et al. (2025)
A IEGPS-CSIC Instituto de Estudios Gallegos Padre Sarmiento (IEGPS-CSIC) Sarymsakova and Martin-rodilla (2025)
A INFOTEC-NLP INFOTEC Centro de Investigación e Innovación en, Tecnologías de la Información y Comunicación, Secretaría de Ciencia, Humanidades, Tecnología e Innovación (SECIHTI) Santos-rodriguez and Graff (2025)
A ITF-NLP Fraunhofer FKIE Kent and Nindel (2025)
A, B JNLP National Institute of Informatics, Japan Advanced Institute of Science and Technology, National Institute of Advanced Industrial Science and Technology Xue et al. (2025)
A JUNLP_Sarika Jadavpur University, Kolkata, India Khatun et al. (2025)
A JellyK University of Information Technology VNU-HCM Le and Thin (2025)
A KDBERT MLDistill none Wang (2025)
A, B LATE-GIL-NLP Universidad Nacional Autonoma de Mexico Vázquez-osorio et al. (2025)
A, C Lazarus NLP University of New South Wales, Lazarus NLP Wongso et al. (2025)
A Lotus PersianGulf University Ranjbar and Baghbani (2025)
A MRS Michigan State University Afshari et al. (2025)
A McGill-NLP Université de Montréal, Mila - Quebec AI Institute, McGill University, Canada CIFAR AI Chair Verma and Adelani (2025)
A NCL-NLP Newcastle University, Newcastle Upon Tyne, England Lu (2025)
A NITK-VITAL National Institute of Technology Karnataka, Surathkal, India Kesanam et al. (2025)
A NLP-Cimat Center for Research in Mathematics (CIMAT), Secretaría de Ciencia, Humanidades, Tecnología e Innovación (SECIHTI) Gómez et al. (2025)
A NLP-DU CSE, DU Sakib et al. (2025)
A, B NLP_goats Sri Sivasubramaniya Nadar (SSN) College of Engineering Vaidyanathan et al. (2025)
A NTA University of Information Technology, Vietnam National University, Ho Chi Minh City Le et al. (2025b)
A NUST Titans National University of Science and Technology, Islamabad Pakistan, Mid Sweden University, Sweden Khan et al. (2025)
A, B NYCU-NLP National Yang Ming Chiao Tung University Xu et al. (2025)
A OPI-DRO-HEL National Information Processing Institute, Warsaw, Poland Karaś and Śpiewak (2025)
A, B, C OZemi Waseda University Takahashi et al. (2025)
A, B PAI none Ruan et al. (2025)
A Pateam PING AN LIFE INSURANCE COMPANY OF CHINA, LTD. Chen et al. (2025b)
A, B Pixel Phantoms SSN college of engineering S et al. (2025a)
A PromotionGo Hubei University Huang and Cui (2025)
A QiMP Technical University of Munich Bogatyreva et al. (2025)
A RSSN Sri Sivasubramaniya Nadar (SSN) College of Engineering V et al. (2025)
A STFXNLP none Murrant et al. (2025)
B Stanford MLab Stanford University Le et al. (2025a)
A, B SyntaxMind none Riad and Ullah (2025)
A, B TILeN none Reyes-magaña et al. (2025)
A Tarbiat Modares Sharif University of Technology, Tarbiat Modares University Bourbour et al. (2025)
A Team A HKBK College of Engineering, Bangalore, India, National Institute of Technology Agartala, Tripura, India Sahil and Jamatia (2025)
A Team KiAmSo University of Tübingen Sharp et al. (2025)
A, B, C Team UBD University of Bucharest Paduraru (2025)
A Team Unibuc - NLP University of Bucharest Creanga et al. (2025)
A TechSSN3 SSN College of Engineering S et al. (2025b)
A, B TeleAI Institute of Artificial Intelligence (TeleAI), China Telecom Corp Ltd Wang et al. (2025)
A Tewodros none Bizuneh et al. (2025)
B Tewodros none Bizuneh et al. (2025)
A Trans-Sent Aliah University, Kolkata, Jadavpur University, Kolkata Sarif et al. (2025)
A, B, C Tue-JMS Universität Tübingen Han et al. (2025)
A, B, C UB_Tel-U Brawijaya University, Telkom University Fatyanosa et al. (2025)
A, B, C UIMP-Aaman none Aman-parveen (2025)
B, C UT-NLP none Safdarian et al. (2025)
A, B, C UncleLM Iran University of Science and Technology Barfi et al. (2025)
A University of Indonesia Universitas Indonesia’ Hanif et al. (2025)
A, C UoB-NLP University of Birmingham Leon et al. (2025)
A VerbaNexAI Universidad Tecnológica de Bolívar Almanza et al. (2025)
A XLM-Muriel Amirkabir University of Technology, Tehran, Iran Hosseinzadeh (2025)
A YNU-NPCC none Yang et al. (2025)
A, B, C YNWA_PZ Iran University of Science and Technology Poulaei et al. (2025)
A, B Zero IIIT Hyderabad Gundam et al. (2025)
A Zero IIIT-Hyderabad Gundam et al. (2025)
A, B, C Zero_Shot Chittagong University of Engineering and Technology Paran et al. (2025)
A, B indiDataMiner Indian Institute of Technology Guwahati, India, NTU Singapore Kumar et al. (2025)
A, B, C maomao none Chen (2025)
B VebanexAI Moreno et al. (2025)
B tinaal Newcastle University, UK Zhu et al. (2025)
A wangkongqiang none Kongqiang (2025)
A xiacui Manchester Metropolitan University Cui (2025)
A zhouyijiang1 University of Yunnan Jiang and Zhang (2025)

### v2

@@
 We present our shared task on text-based emotion detection, covering more than 30 languages from seven distinct language families.
-These languages are predominantly low-resource and spoken across various continents.
-The data instances are multi-labeled into six emotional classes, with additional datasets in 11 languages annotated for emotion intensity.
-Participants were asked to predict labels in three tracks: (a) emotion labels in monolingual settings, (b) emotion intensity scores, and (c) emotion labels in cross-lingual settings.
+These languages are predominantly low-resource and are spoken across various continents.
+The data instances are multi-labeled with six emotional classes, with additional datasets in 11 languages annotated for emotion intensity.
+Participants were asked to predict labels in three tracks: (a) multilabel emotion detection, (b) emotion intensity score detection, and (c) cross-lingual emotion detection.
 The task attracted over 700 participants.
 We received final submissions from more than 200 teams and 93 system description papers.
-We report baseline results, as well as findings on the best-performing systems, the most common approaches, and the most effective methods across various tracks and languages.
+We report baseline results, along with findings on the best-performing systems, the most common approaches, and the most effective methods across different tracks and languages.
 The datasets for this task are publicly available.
-[^1] The dataset is available at SemEval2024-task 11.
+[^1] The dataset is available at SemEval2025 Task 11.
 […]
@@
 (2023).
-In this work, we use emotion recognition to refer to perceived emotions—i.e., the emotion most people believe the speaker might have felt based on a sentence or short text snippet.
+In this work, we use emotion recognition to refer to perceived emotions, i.e., the emotion most people believe the speaker might have felt based on a sentence or short text snippet.
 Figure 1: Languages in the three tracks (A, B, and C) of SemEval 2025 Task 11.
@@
 Additionally, eleven datasets include four emotional intensity levels ranging from 0 to 3 (i.e., absence of emotion to high intensity).
-The task consists of three tracks: (a) multi-label emotion detection, (b) emotion intensity detection, and (c) cross-lingual emotion detection.
+The task consists of three tracks: (a) multilabel emotion detection, (b) emotion intensity detection, and (c) cross-lingual emotion detection.
 The languages for each track are listed in Figure 1.
@@
 Most teams participated in multiple languages, averaging 11 languages per team.
-Our task was the most popular competition on CodaBench in 2024.
-[^3] Codabench Newsletter 2024 All task details and resources are available on the task’s GitHub page.
+Our task was the most popular competition on CodaBench in 2024.333Codabench Newsletter 2024 All task details and resources are available on the task’s GitHub page.
 [^4] https://github.com/emotion-analysis-project/SemEval2025-Task11
@@
 While multilingual datasets (e.g., (Öhman et al., 2020) and XLM-EMO (Bianchi et al., 2022)) exist, they do not fully capture cultural nuances in emotional expressions due to their reliance on translated data (e.g., XLM-EMO), as emotions are highly contextualized and culture-specific (Havaldar et al., 2023; Mohamed et al., 2024; Hershcovich et al., 2022).
-Furthermore, most datasets are single-labeled, and to the best of our knowledge, there are no multilingual resources that capture overlapping emotions and their intensity across various languages.
-Additionally, most prior emotion recognition shared tasks have focused on high-resource languages such as English, Spanish, German, and Arabic Strapparava and Mihalcea (2007); Mohammad et al.
+Furthermore, most datasets are single-labeled, and to the best of our knowledge, there are no multilingual resources that capture simultaneous emotions and their intensity across various languages.
+Additionally, most prior emotion recognition shared tasks have focused on high-resource languages such as English, Spanish, German, and Arabic Strapparava and Mihalcea (2007); Mohammad and Bravo-Marquez (2017); Mohammad et al.
 (2018b); Chatterjee et al.
 (2019).
-In contrast, this shared task covers more than 30 languages, including many low-resource languages.
+In contrast, this shared task covers more than 30 languages, including several low-resource languages.
 3 Data
@@
 Similarly, the afr dataset includes sentences from speeches and talks.
-  • Literary texts: We translated the novel *La Grande Maison* (The Big House) by Mohammed Dib  [^5] https://en.wikipedia.org/wiki/La_Grande_Maison from French to Algerian Arabic (arq) and manually post-processed the translation to generate sentences to be annotated by native speakers.
-  • News data: Although we prefer emotionally rich social media data from different platforms, when such data is scarce, we annotate news data and headlines in some African languages (e.g., yor, hau, and vmw).
+  • Literary texts: The language lead manually translated the novel La Grande Maison (The Big House) by Mohammed Dib [^5] https://en.wikipedia.org/wiki/La_Grande_Maison from French into Algerian Arabic (arq), and post-processed the translation to generate sentences for annotation by native speakers.
+Note that the translator is bilingual and a native speaker of Algerian Arabic.
+  • News data: Although we prefer emotionally rich social media data from different platforms, when such data is scarce, we annotated news data and headlines in some African languages (e.g., yor, hau, and vmw).
   • Human-written and machine-generated data: We created a dataset from scratch for Hindi (hin) and Marathi (mar).
-We ask annotators to generate emotive sentences on a given topic (e.g., family), automatically translate a small section of the Hindi dataset to Marathi, and have native speakers manually fix translation errors.
-Finally, we augment both datasets with a few hundred quality-approved instances generated by ChatGPT.
+Annotators were asked to come up with emotive sentences on a given topic (e.g., family).
+A small portion of the Hindi dataset was automatically translated into Marathi and manually corrected by native speakers to fix translation errors.
+Finally, we augmented both datasets with a few hundred quality-approved instances generated by ChatGPT.
+Note that these constitute less than 1% of the total number of data instances.
 3.2 Data Annotation
@@
     The average score exceeds a predefined threshold $T$. We set $T$ to $0.5$.
-Once the labels for perceived emotions are assigned, the final intensity scores for Track B were determined by averaging the selected intensity values and rounding up to the nearest whole number.
-We only assign intensity scores for datasets where most instances were annotated by at least five annotators to ensure robustness.
-Table 1 shows the total number of instances in each dataset and the number of instances in the training, development, and test splits for all the languages.
+Once the perceived emotion labels are assigned, the final intensity scores for Track B are determined by averaging the selected intensity values and rounding up to the nearest whole number.
+Intensity scores are assigned only for datasets in which most instances were annotated by at least five annotators to ensure robustness.
+Table 1 shows the total number of instances in each dataset, as well as the number of instances in the training, development, and test splits for all languages.
 3.3 Annotators’ Reliability
@@
 Track B: Emotion Intensity Detection
-Given a text and the six emotion classes (i.e., joy, sadness, fear, anger, surprise, and disgust), participants had to predict whether the intensity of each possible emotion was equal to 0 (no emotion), 1 (low), 2 (medium), or 3 (high).
+Given a text and six emotion classes (i.e., joy, sadness, fear, anger, surprise, and disgust), participants were required to predict whether the intensity of each emotion was 0 (no emotion), 1 (low), 2 (medium), or 3 (high).
+Note that Track B does not include all languages, as intensity scores were only released for datasets with at least five annotators per instance to ensure more robust and reliable labels.
 Track C: Cross-lingual Emotion Detection
-Similar to Track A, participants had to predict the presence or absence of each perceived emotion, but without using any training data in the target language.
-Instead, participants were allowed to use labeled dataset(s) from at least one other language.
+Similar to Track A, participants were required to predict the presence or absence of each perceived emotion, but without using any training data in the target language.
+Instead, they were permitted to use labeled dataset(s) from at least one other language.
 For instance, one could use German data for training when testing on English.
+This track focuses on cross-lingual transfer and explores how data from various languages can support emotion detection in low-resource settings, as well as the ability of models to generalise across domains.
 4.2 Task Organisation
@@
 The list includes 33 countries: Argentina, Australia, Bangladesh, Brazil, Canada, Colombia, Egypt, Ethiopia, Finland, Germany, Greece, India, Indonesia, Iran, Japan, Jordan, Mexico, Morocco, the Netherlands, Nigeria, Pakistan, Poland, Romania, Russia, South Africa, Spain, Sweden, Taiwan, the UAE, the UK, the USA, and Vietnam.
+Lang Team Score Lang Team Score Lang Team Score Lang Team Score
+afr pai 0.699 amh chinchunmei 0.773 arq pai 0.669 ary pai 0.629
+maomao 0.687 nust titans 0.714 jnlp 0.641 jnlp 0.609
+$R_{\text{baseline}}$ 0.371 $R_{\text{baseline}}$ 0.638 $R_{\text{baseline}}$ 0.414 $R_{\text{baseline}}$ 0.472
+$M_{\text{baseline}}$ 0.257 $M_{\text{baseline}}$ 0.295 $M_{\text{baseline}}$ 0.445 $M_{\text{baseline}}$ 0.247
+chn pai 0.709 deu pai 0.740 eng pai 0.823 esp pai 0.849
+teleai 0.682 heimerdinger 0.706 nycu-nlp 0.823 heimerdinger 0.838
+$R_{\text{baseline}}$ 0.531 $R_{\text{baseline}}$ 0.642 $R_{\text{baseline}}$ 0.708 $R_{\text{baseline}}$ 0.774
+$M_{\text{baseline}}$ 0.278 $M_{\text{baseline}}$ 0.449 $M_{\text{baseline}}$ 0.367 $M_{\text{baseline}}$ 0.312
+hau pai 0.751 hin jnlp 0.926 ibo pai 0.600 kin pai 0.657
+empaths 0.695 pai 0.920 late-gil-nlp 0.563 mcgill-nlp 0.590
+$R_{\text{baseline}}$ 0.596 $R_{\text{baseline}}$ 0.855 $R_{\text{baseline}}$ 0.479 $R_{\text{baseline}}$ 0.463
+$M_{\text{baseline}}$ 0.312 $M_{\text{baseline}}$ 0.246 $M_{\text{baseline}}$ 0.236 $M_{\text{baseline}}$ 0.218
+mar pai 0.884 orm tewodros 0.616 pcm pai 0.674 ptbr pai 0.683
+indidataminer 0.883 late-gil-nlp 0.592 jnlp 0.634 heimerdinger 0.625
+$R_{\text{baseline}}$ 0.822 $R_{\text{baseline}}$ 0.126 $R_{\text{baseline}}$ 0.555 $R_{\text{baseline}}$ 0.426
+$M_{\text{baseline}}$ 0.264 $M_{\text{baseline}}$ 0.232 $M_{\text{baseline}}$ 0.357 $M_{\text{baseline}}$ 0.243
+ptmz pai 0.548 ron pai 0.794 rus heimerdinger 0.901 som pai 0.577
+heimerdinger 0.507 jnlp 0.779 jnlp 0.891 empaths 0.508
+$R_{\text{baseline}}$ 0.459 $R_{\text{baseline}}$ 0.762 $R_{\text{baseline}}$ 0.838 $R_{\text{baseline}}$ 0.459
+$M_{\text{baseline}}$ 0.163 $M_{\text{baseline}}$ 0.461 $M_{\text{baseline}}$ 0.262 $M_{\text{baseline}}$ 0.198
+sun lazarus nlp 0.550 swa empaths 0.386 swe pai 0.626 tat pai 0.846
+pai 0.541 pai 0.385 jnlp 0.619 tue-jms 0.797
+$R_{\text{baseline}}$ 0.373 $R_{\text{baseline}}$ 0.227 $R_{\text{baseline}}$ 0.520 $R_{\text{baseline}}$ 0.539
+$M_{\text{baseline}}$ 0.334 $M_{\text{baseline}}$ 0.179 $M_{\text{baseline}}$ 0.264 $M_{\text{baseline}}$ 0.246
+tir nta 0.591 ukr pai 0.726 vmw team unibuc 0.325 yor pai 0.461
+late-gil-nlp 0.587 csiro-lt 0.664 pai 0.255 heimerdinger 0.392
+$R_{\text{baseline}}$ 0.463 $R_{\text{baseline}}$ 0.535 $R_{\text{baseline}}$ 0.121 $R_{\text{baseline}}$ 0.092
+$M_{\text{baseline}}$ 0.253 $M_{\text{baseline}}$ 0.157 $M_{\text{baseline}}$ 0.163 $M_{\text{baseline}}$ 0.165
+Table 2: Average macro-F1 scores for our baselines ($M_{baseline}$ and $R_{baseline}$, referring to the Majority Vote and RoBERTa baselines, respectively) and the top two performing systems in Track A (shown in bold) for each language.
 We used Codabench as the competition platform and released pilot datasets before the start of the shared task to help participants better understand the task (i.e., the datasets, the languages involved, and the labels).
-We provided participants with a starter kit on GitHub, resources for beginners, and organised a Q&A session along with a writing tutorial.
+We provided participants with a starter kit on GitHub, resources for beginners, and organised a Q&A session along with a writing tutorial for junior researchers.
 Our participants were based in different parts of the world, as shown in Figure 2, with many coming from underrepresented regions.
@@
 5 Participating Systems and Results
-Lang Team Score Lang Team Score Lang Team Score Lang Team Score
-afr pai 0.699 amh chinchunmei 0.773 arq pai 0.669 ary pai 0.629
-maomao 0.687 nust titans 0.714 jnlp 0.641 jnlp 0.609
-$R_{\text{baseline}}$ 0.371 $R_{\text{baseline}}$ 0.638 $R_{\text{baseline}}$ 0.414 $R_{\text{baseline}}$ 0.472
-$M_{\text{baseline}}$ 0.257 $M_{\text{baseline}}$ 0.295 $M_{\text{baseline}}$ 0.445 $M_{\text{baseline}}$ 0.247
-chn pai 0.709 deu pai 0.740 eng pai 0.823 esp pai 0.849
-teleai 0.682 heimerdinger 0.706 nycu-nlp 0.823 heimerdinger 0.838
-$R_{\text{baseline}}$ 0.531 $R_{\text{baseline}}$ 0.642 $R_{\text{baseline}}$ 0.708 $R_{\text{baseline}}$ 0.774
-$M_{\text{baseline}}$ 0.278 $M_{\text{baseline}}$ 0.449 $M_{\text{baseline}}$ 0.367 $M_{\text{baseline}}$ 0.312
-hau pai 0.751 hin jnlp 0.926 ibo pai 0.600 kin pai 0.657
-empaths 0.695 pai 0.920 late-gil-nlp 0.563 mcgill-nlp 0.590
-$R_{\text{baseline}}$ 0.596 $R_{\text{baseline}}$ 0.855 $R_{\text{baseline}}$ 0.479 $R_{\text{baseline}}$ 0.463
-$M_{\text{baseline}}$ 0.312 $M_{\text{baseline}}$ 0.246 $M_{\text{baseline}}$ 0.236 $M_{\text{baseline}}$ 0.218
-mar pai 0.884 orm tewodros 0.616 pcm pai 0.674 ptbr pai 0.683
-indidataminer 0.883 late-gil-nlp 0.592 jnlp 0.634 heimerdinger 0.625
-$R_{\text{baseline}}$ 0.822 $R_{\text{baseline}}$ 0.126 $R_{\text{baseline}}$ 0.555 $R_{\text{baseline}}$ 0.426
-$M_{\text{baseline}}$ 0.264 $M_{\text{baseline}}$ 0.232 $M_{\text{baseline}}$ 0.357 $M_{\text{baseline}}$ 0.2425
-ptmz pai 0.548 ron pai 0.794 rus heimerdinger 0.901 som pai 0.577
-heimerdinger 0.507 jnlp 0.779 jnlp 0.891 empaths 0.508
-$R_{\text{baseline}}$ 0.459 $R_{\text{baseline}}$ 0.762 $R_{\text{baseline}}$ 0.838 $R_{\text{baseline}}$ 0.459
-$M_{\text{baseline}}$ 0.163 $M_{\text{baseline}}$ 0.461 $M_{\text{baseline}}$ 0.262 $M_{\text{baseline}}$ 0.198
-sun lazarus nlp 0.550 swa empaths 0.386 swe pai 0.626 tat pai 0.846
-pai 0.541 pai 0.385 jnlp 0.619 tue-jms 0.797
-$R_{\text{baseline}}$ 0.373 $R_{\text{baseline}}$ 0.227 $R_{\text{baseline}}$ 0.520 $R_{\text{baseline}}$ 0.539
-$M_{\text{baseline}}$ 0.334 $M_{\text{baseline}}$ 0.179 $M_{\text{baseline}}$ 0.264 $M_{\text{baseline}}$ 0.246
-Table 2: Average macro-F1 scores for our baselines ($M_{baseline}$ and $R_{baseline}$ referring to Majority Vote and RoBERTa baselines, respectively) and the top two performing systems for each language in Track A (in bold).
 5.1 Overview
@@
 They consistently rank as the top approach in Track A for 20 out of 28 languages.
-For their system, they combine several base models (ChatGPT-4o, DeepSeek-V3, Gemma-9b, Qwen-2.5-32b, Mistral-Small-24B) using multiple ensemble techniques (neural networks, XGBoost, LightGBM, linear regression, weighted voting).
-They fine-tune the Gemma and Qwen models using AdaLoRA.
+For their system, they combine several base models (ChatGPT-4o OpenAI (2024), DeepSeek-V3 DeepSeek-AI et al.
+(2025), Gemma-9b Team et al.
+(2024), Qwen-2.5-32b Yang et al.
+(2024), Mistral-Small-24B Jiang et al.
+(2024)) using multiple ensemble techniques (neural networks, XGBoost, LightGBM, linear regression, weighted voting).
+They fine-tune Gemma-9b and Qwen-2.5-32b using AdaLoRA.
 For prompting the LLMs, they used an iterative prompt-optimisation technique that generates prompt variations.
@@
 Their samples are randomly selected from the task dataset (no external augmentation).
-They use LLaMa3-Instruct-8B for their fine-tuning.
+They use LLaMa3-Instruct-8B AI@Meta (2024) for their fine-tuning.
 5.2.2 Takeaways
-Most of the teams that rank well on Track A experiment with essentially two methodologies: 1) fine-tuning BERT-based models such as DeBERTa, mBERT, and XLM-R; and/or 2) instruction-fine-tuning using LoRa methodologies in combination with prompt design and data augmentation techniques on LLMs (ChatGPT-4o, DeepSeek-V3, Gemma-9b, Qwen-2.5-32b, Mistral-Small-24B).
-For instance, Team Telai (3rd in Chinese), Team Empaths (2nd in Hausa and Somali, and 1st in Swahili), Team NYCU-NLP (2nd in English), Team JNLP (1st in Hindi and among the best four across 10 languages), Team Unibouc (1st in Emakhuwa), Team Heiderdinger (2nd in Mozambican Portuguese, German, Spanish, Brazilian Portuguese, and Yoruba; 1st in Russian), and Team Maomao (2nd in Afrikaans).
+Most of the teams that rank well on Track A experiment with essentially two methodologies: 1) fine-tuning BERT-based models such as DeBERTa Siino (2024), mBERT Dolev (2023), and XLM-R Conneau et al.
+(2020); and/or 2) instruction-fine-tuning using LoRa methodologies in combination with prompt design and data augmentation techniques on LLMs (ChatGPT-4o, DeepSeek-V3, Gemma-9b, Qwen-2.5-32b, Mistral-Small-24B).
+For instance, Team Telai (3rd in Chinese), Team Empaths (2nd in Hausa and Somali, and 1st in Swahili), Team NYCU-NLP (2nd in English), Team JNLP (1st in Hindi and among the best four across 10 languages), Team Unibouc (1st in Emakhuwa), Team Heiderdinger (2nd in Mozambican Portuguese, German, Spanish, Brazilian Portuguese, and Yorùbá; 1st in Russian), and Team Maomao (2nd in Afrikaans).
 Few teams focus on only a subset of languages or explore language-related knowledge in their methodology.
-For instance, Team Lazarus NLP redefined and reformulated the multilabel classification into multiple binary tasks to expand training samples.
+For instance, Team Lazarus NLP redefined and reformulated the multi-label classification into multiple binary tasks to expand training samples.
 They also explored how knowledge could potentially be transferred between Indonesian languages.
-All the top 10 teams performed much better than our baseline model, with an improvement that is more notable in a few low-resource languages, such as Oromo, where Team Tewodros obtained an average macro-F1 score of 0.616 compared to a baseline of 0.126.
-The same was observed for Yoruba, where Team Pai scored 0.461 compared to a baseline score of 0.092.
+All the top 10 teams performed significantly better than our baseline model, with an improvement that is more notable in a few low-resource languages, such as Oromo, where Team Tewodros obtained an average macro-F1 score of 0.616 compared to a baseline of 0.126.
+The same was observed for Yorùbá, where Team Pai scored 0.461 compared to a baseline score of 0.092.
 5.3 Track B: Emotion Intensity Detection
@@
 teleai 0.708 teleai 0.743
-$R_{\text{baseline}}$ 0.405 0.562
+$R_{\text{baseline}}$ 0.405 $R_{\text{baseline}}$ 0.562
 $M_{\text{baseline}}$ 0.000 $M_{\text{baseline}}$ 0.016
@@
 $M_{\text{baseline}}$0.01
-Table 3: Pearson correlation scores for our baselines (Majority $M_{baseline}$ and RoBERTa $R_{baseline}$) and the top 2 performing systems for each language in Track B (in bold).
+Table 3: Pearson correlation scores for our baselines (Majority: $M_{baseline}$ and RoBERTa: $R_{baseline}$) and the top two performing systems in Track B (shown in bold) for each language.
+Lang Team Score Lang Team Score Lang Team Score Lang Team Score
+afr maomao 0.705 amh deepwave 0.661 arq deepwave 0.588 ary deepwave 0.632
+deepwave 0.574 uob-nlp 0.627 maomao 0.584 maomao 0.565
+$R_{\text{baseline}}$ 0.350 $R_{\text{baseline}}$ 0.487 $R_{\text{baseline}}$ 0.338 $R_{\text{baseline}}$ 0.355
+$M_{\text{baseline}}$ 0.257 $M_{\text{baseline}}$ 0.295 $M_{\text{baseline}}$ 0.445 $M_{\text{baseline}}$ 0.247
+chn deepwave 0.689 deu deepwave 0.727 eng deepwave 0.797 esp deepwave 0.831
+maomao 0.622 gt-nlp 0.687 maomao 0.755 maomao 0.806
+$R_{\text{baseline}}$ 0.246 $R_{\text{baseline}}$ 0.468 $R_{\text{baseline}}$ 0.375 $R_{\text{baseline}}$ 0.574
+$M_{\text{baseline}}$ 0.278 $M_{\text{baseline}}$ 0.319 $M_{\text{baseline}}$ 0.449 $M_{\text{baseline}}$ 0.367
+hau deepwave 0.709 hin deepwave 0.919 ibo deepwave 0.605 ind maomao 0.672
+uob-nlp 0.627 maomao 0.896 uob-nlp 0.484 lazarus nlp 0.641
+$R_{\text{baseline}}$ 0.320 $R_{\text{baseline}}$ 0.138 $R_{\text{baseline}}$ 0.075 $R_{\text{baseline}}$ 0.376
+$M_{\text{baseline}}$ 0.312 $M_{\text{baseline}}$ 0.264 $M_{\text{baseline}}$ 0.236 $M_{\text{baseline}}$ 0.254
+jav heimerdinger 0.439 kin deepwave 0.508 mar deepwave 0.903 orm deepwave 0.542
+lazarus nlp 0.438 uob-nlp 0.466 maomao 0.863 uob-nlp 0.491
+$R_{\text{baseline}}$ 0.464 $R_{\text{baseline}}$ 0.184 $R_{\text{baseline}}$ 0.772 $R_{\text{baseline}}$ 0.262
+$M_{\text{baseline}}$ 0.204 $M_{\text{baseline}}$ 0.218 $M_{\text{baseline}}$ 0.264 $M_{\text{baseline}}$ 0.232
+pcm deepwave 0.674 ptbr deepwave 0.629 ptmz deepwave 0.555 ron deepwave 0.767
+maomao 0.562 maomao 0.617 maomao 0.495 maomao 0.747
+$R_{\text{baseline}}$ 0.010 $R_{\text{baseline}}$ 0.418 $R_{\text{baseline}}$ 0.297 $R_{\text{baseline}}$ 0.762
+$M_{\text{baseline}}$ 0.357 $M_{\text{baseline}}$ 0.243 $M_{\text{baseline}}$ 0.163 $M_{\text{baseline}}$ 0.652
+rus deepwave 0.906 som maomao 0.488 sun deepwave 0.467 swa maomao 0.381
+maomao 0.852 deepwave 0.488 maomao 0.464 deepwave 0.355
+$R_{\text{baseline}}$ 0.704 $R_{\text{baseline}}$ 0.273 $R_{\text{baseline}}$ 0.194 $R_{\text{baseline}}$ 0.190
+$M_{\text{baseline}}$ 0.262 $M_{\text{baseline}}$ 0.198 $M_{\text{baseline}}$ 0.334 $M_{\text{baseline}}$ 0.179
+swe deepwave 0.645 tat deepwave 0.789 tir deepwave 0.505 ukr deepwave 0.702
+maomao 0.578 maomao 0.697 uob-nlp 0.445 maomao 0.623
+$R_{\text{baseline}}$ 0.512 $R_{\text{baseline}}$ 0.445 $R_{\text{baseline}}$ 0.339 $R_{\text{baseline}}$ 0.496
+$M_{\text{baseline}}$ 0.264 $M_{\text{baseline}}$ 0.246 $M_{\text{baseline}}$ 0.253 $M_{\text{baseline}}$ 0.157
+vmw deepwave 0.210 xho maomao 0.443 yor maomao 0.360 zul maomao 0.397
+ozemi 0.193 ozemi 0.315 deepwave 0.342 heimerdinger 0.226
+$R_{\text{baseline}}$ 0.052 $R_{\text{baseline}}$ 0.127 $R_{\text{baseline}}$ 0.053 $R_{\text{baseline}}$ 0.153
+$M_{\text{baseline}}$ 0.162 $M_{\text{baseline}}$ 0.115 $M_{\text{baseline}}$ 0.165 $M_{\text{baseline}}$ 0.109
+Table 4: Average macro-F1 scores for our baselines ($M_{baseline}$ and $R_{baseline}$, referring to the Majority Vote and RoBERTa baselines, respectively) and the top two performing systems in Track C (shown in bold) for each language.
 5.3.1 Best-Performing Systems
@@
 5.3.2 Takeaways
-Teams Deepwave, Teleai, and JNLP also ranked highly across various languages using similar prompt engineering approaches as in Track A.
+Teams Deepwave, Teleai, and JNLP also ranked highly across various languages using prompt engineering approaches similar to those in Track A.
 Additionally, Team NYCU-NLP ranked second in English by aggregating instruction-tuned small language models.
-The different teams outperformed our RoBERTa baseline, which achieved medium Pearson correlation coefficient scores overall, with some extremely low scores in Algerian Arabic, Hausa, Ukrainian, and even Brazilian Portuguese.
-Overall, we observe that the teams adopted approaches similar to Track A’s, with minor adjustments in the prompts.
-Furthermore, we observe that even the best-performing teams achieved a Pearson correlation coefficient score of no more than 0.65 on Algerian Arabic, likely due to the novelty of the dataset.
+All these teams outperformed our RoBERTa baseline, which achieved moderate Pearson correlation coefficient scores overall, but performed poorly in languages such as Algerian Arabic, Hausa, Ukrainian, and even Brazilian Portuguese -highlighting the difficulty of the task.
+Overall, we observe that most teams adopted approaches similar to those used in Track A, with only minor adjustments to the prompts.
+Notably, even the best-performing teams achieved a Pearson correlation coefficient of no more than 0.65 on Algerian Arabic, likely due to the novelty and complexity of the dataset.
 5.4 Track C: Cross-lingual Emotion Detection
-Lang Team Score Lang Team Score Lang Team Score Lang Team Score
-afr maomao 0.705 amh deepwave 0.661 arq deepwave 0.588 ary deepwave 0.632
-deepwave 0.574 uob-nlp 0.627 maomao 0.584 maomao 0.565
-$R_{\text{baseline}}$ 0.350 $R_{\text{baseline}}$ 0.487 $R_{\text{baseline}}$ 0.338 $R_{\text{baseline}}$ 0.355
-$M_{\text{baseline}}$ 0.257 $M_{\text{baseline}}$ 0.295 $M_{\text{baseline}}$ 0.445 $M_{\text{baseline}}$ 0.247
-chn deepwave 0.689 deu deepwave 0.727 eng deepwave 0.797 esp deepwave 0.831
-maomao 0.622 gt-nlp 0.687 maomao 0.755 maomao 0.806
-$R_{\text{baseline}}$ 0.246 $R_{\text{baseline}}$ 0.468 $R_{\text{baseline}}$ 0.375 $R_{\text{baseline}}$ 0.574
-$M_{\text{baseline}}$ 0.278 $M_{\text{baseline}}$ 0.319 $M_{\text{baseline}}$ 0.449 $M_{\text{baseline}}$ 0.367
-hau deepwave 0.709 hin deepwave 0.919 ibo deepwave 0.605 ind maomao 0.672
-uob-nlp 0.627 maomao 0.896 uob-nlp 0.484 lazarus nlp 0.641
-$R_{\text{baseline}}$ 0.320 $R_{\text{baseline}}$ 0.138 $R_{\text{baseline}}$ 0.075 $R_{\text{baseline}}$ 0.376
-$M_{\text{baseline}}$ 0.312 $M_{\text{baseline}}$ 0.264 $M_{\text{baseline}}$ 0.236 $M_{\text{baseline}}$
-jav heimerdinger 0.439 kin deepwave 0.508 mar deepwave 0.903 orm deepwave 0.542
-lazarus nlp 0.438 uob-nlp 0.466 maomao 0.863 uob-nlp 0.491
-$R_{\text{baseline}}$ 0.464 $R_{\text{baseline}}$ 0.184 $R_{\text{baseline}}$ 0.772 $R_{\text{baseline}}$ 0.262
-$M_{\text{baseline}}$ $M_{\text{baseline}}$ 0.218 $M_{\text{baseline}}$ 0.264 $M_{\text{baseline}}$ 0.232
-pcm deepwave 0.674 ptbr deepwave 0.629 ptmz deepwave 0.555 ron deepwave 0.767
-maomao 0.562 maomao 0.617 maomao 0.495 maomao 0.747
-$R_{\text{baseline}}$ 0.010 $R_{\text{baseline}}$ 0.418 $R_{\text{baseline}}$ 0.297 $R_{\text{baseline}}$ 0.762
-$M_{\text{baseline}}$ 0.357 $M_{\text{baseline}}$ 0.243 $M_{\text{baseline}}$ 0.163 $M_{\text{baseline}}$
-rus deepwave 0.906 som maomao 0.488 sun deepwave 0.467 swa maomao 0.381
-maomao 0.852 deepwave 0.488 maomao 0.464 deepwave 0.355
-$R_{\text{baseline}}$ 0.704 $R_{\text{baseline}}$ 0.273 $R_{\text{baseline}}$ 0.194 $R_{\text{baseline}}$ 0.190
-$M_{\text{baseline}}$ 0.262 $M_{\text{baseline}}$ 0.198 $M_{\text{baseline}}$ 0.334 $M_{\text{baseline}}$ 0.179
-swe deepwave 0.645 tat deepwave 0.789 tir deepwave 0.505 ukr deepwave 0.702
-maomao 0.578 maomao 0.697 uob-nlp 0.445 maomao 0.623
-$R_{\text{baseline}}$ 0.512 $R_{\text{baseline}}$ 0.445 $R_{\text{baseline}}$ 0.339 $R_{\text{baseline}}$ 0.496
-$M_{\text{baseline}}$ 0.264 $M_{\text{baseline}}$ 0.246 $M_{\text{baseline}}$ 0.253 $M_{\text{baseline}}$ 0.157
-vmw deepwave 0.210 xho maomao 0.443 yor maomao 0.360 zul maomao 0.397
-ozemi 0.193 ozemi 0.315 deepwave 0.342 heimerdinger 0.226
-$R_{\text{baseline}}$ 0.052 $R_{\text{baseline}}$ 0.127 $R_{\text{baseline}}$ 0.053 $R_{\text{baseline}}$ 0.153
-$M_{\text{baseline}}$ 0.162 $M_{\text{baseline}}$ $M_{\text{baseline}}$ 0.165 $M_{\text{baseline}}$
-Table 4: Average macro-F1 scores for our baselines ($M_{baseline}$ and $R_{baseline}$ referring to Majority Vote and RoBERTa baselines, respectively) and the top 2 performing systems for each language in Track C (in bold).
+Figure 3: Top LLMs used by participants across tracks (A, B, and C).
 5.4.1 Best-Performing Systems
 Team deepwave
-Team Deepwave fine-tuned Google Gemma-2 using tailored data augmentation and Chain-of-Thought (CoT) prompting.
+Team Deepwave fine-tuned Google Gemma-2 Team et al.
+(2024) using tailored data augmentation and Chain-of-Thought (CoT) prompting.
 They decomposed the task into two sub-tasks: (1) sentiment keyword identification and (2) sentiment polarity recognition.
@@
 Team maomao experimented with different setups for fine-tuning LLMs.
-The base models used were Qwen2.5-7B-Instruct, GPT-0544o Mini2, and LLaMA-3.2-3B-Instruct (Touvron et al., 2023).
-They applied Direct Preference Optimization to refine their model—a technique that selects high-quality instances from lower-quality ones within a dataset.
+The base models used were Qwen2.5-7B-Instruct, GPT-0544o Mini2, and LLaMA-3.2-3B-Instruct (AI@Meta, 2024).
+They applied Direct Preference Optimisation to refine their model -a technique that selects high-quality instances from lower-quality ones within a dataset.
 After this step, they retrained the model using the refined dataset.
 They also explored random sampling and retrieval-augmented generation (RAG) methods for training, primarily on DeepSeek-V3, Qwen-Max5093, and Grok-V2.
-Figure 3: Top LLMs used by participants across tracks (A, B, and C).
 5.4.2 Takeaways
-Other top-performing systems include Team Ozemi, who fine-tuned a multilingual BERT model and applied machine translation to enhance performance across 35 languages.
+Other top-performing systems include Team Ozemi, who fine-tuned a multilingual BERT model and applied machine translation to enhance performance across all languages.
 They used the Synthetic Minority Oversampling Technique (SMOTE) with TF-IDF to address class imbalance in Russian.
-It appears they translated all datasets into a common language using Google Translate before processing, except for Nigerian Pidgin and Emakhua, where they used a multilingual BERT model for translation.
+They translated all the datasets into a common language using Google Translate before processing -except for Nigerian Pidgin and Emakhua, where they used a multilingual BERT model for translation.
 They also leveraged two Kaggle competition datasets for data augmentation.
@@
 They employed in-context learning with multilingual examples from high-resource languages such as English and Spanish.
-Overall, the different teams outperformed our baselines.
-However, we observe that the average scores for this track are lower, particularly in low-resource languages such as Emakhuwa, as participants were presented with an additional challenge.
+Overall, the participating teams outperformed our baselines.
+However, the average scores for this track are notably lower, particularly in low-resource languages, due to the additional challenges posed by limited data and resources.
+As shown in Table 4, there are significant performance gaps -even the top systems did not achieve an F-score higher than 0.50 in languages such as Javanese, Somali, Sundanese, Xhosa, Yorùbá, isiZulu, and Emakhuwa (where the top system achieved an F-score of only 0.21).
 6 Discussion
@@
 For prompting, few-shot, zero-shot, and chain-of-thought prompting were the most frequently used techniques.
-Many participants also experimented with traditional transformer-based models, particularly XLM-RoBERTa, mBERT, DeBERTa, and IndicBERT (see Figure 3 and Figure 4 in the Appendix).
+Many participants also experimented with traditional transformer-based models, particularly XLM-RoBERTa, mBERT, DeBERTa, and IndicBERT Kakwani et al.
+(2020) (see Figure 3 and Figure 4 in the Appendix).
 Best Performing Systems
-The results from the top-performing submissions suggest that while LLMs achieve strong overall performance, their effectiveness heavily depends on prompt engineering techniques and wording.
+The results from the top-performing submissions suggest that while LLMs achieve strong overall performance, their effectiveness is heavily dependent on prompt engineering techniques and wording.
 Additionally, performance varies significantly by language.
 Across all tracks, LLM-based approaches and the best-performing systems consistently yielded better results for high-resource languages such as English and Russian.
-Conversely, performance dropped notably when tested on low-resource languages such as Swahili and Emakhuwa.
-Furthermore, most teams did not incorporate additional datasets to enhance performance (see Appendix) as few-shot and zero-shot approaches proved highly effective.
+In contrast, performance dropped notably when tested on low-resource languages such as Swahili and Emakhuwa.
+Furthermore, most teams did not incorporate additional datasets to enhance performance (see Appendix), as few-shot and zero-shot approaches proved highly effective.
 7 Conclusion
@@
 The submitted systems were ranked based on macro F1-scores for Tracks A and C, comparing predicted labels to gold labels, and based on the ranking of predicted intensity scores for Track B.
-We summarised the reported results, highlighting the best-performing and most innovative methods.
+We summarised the reported results, discussing the best-performing and most innovative methods.
 Overall, performance varied significantly across languages.
-Limitations
+Our results highlight that emotion recognition remains an open challenge, particularly for under-served languages and in low-resource settings.
+8 Limitations
 Emotions are subjective and subtle, and they are expressed and perceived differently.
 We do not claim that our datasets capture the true emotions of the speakers, fully represent language use across the 32 languages, or cover all possible emotions.
-We discuss these limitations extensively in Ethical Considerations.
+We discuss the ethical considerations extensively in Section 9.
 We acknowledge the limited data sources available for some low-resource languages.
@@
 However, they serve as a valuable starting point for research in this area.
-Ethical Considerations
+9 Ethical Considerations
 Emotion perception and expression are subjective and nuanced, as they are influenced by various factors (e.g., cultural background, social group, personal experiences, and social context).